### v3

# Characterizing the minimax rate of nonparametric regression under bounded convex constraints

Abstract
We quantify the minimax rate for a nonparametric regression model over a convex function class $\mathcal{F}$ with bounded diameter. We obtain a minimax rate of ${\varepsilon^{\ast}}^{2}\wedge\mathrm{diam}(\mathcal{F})^{2}$ where
$\varepsilon^{\ast}=\sup\{\varepsilon>0:n\varepsilon^{2}\leq\log M_{\mathcal{F}}^{\operatorname{loc}}(\varepsilon,c)\},$
where $M_{\mathcal{F}}^{\operatorname{loc}}(\cdot,c)$ is the local metric entropy of $\mathcal{F}$ and our loss function is the squared population $L_{2}$ distance over our input space $\mathcal{X}$. In contrast to classical works on the topic (cf. yang1999information), our results do not require functions in $\mathcal{F}$ to be uniformly bounded in sup-norm. In addition, we prove that our estimator is adaptive to the true point, and to the best of our knowledge this is the first such estimator in this general setting. This work builds on the Gaussian sequence framework of neykov2022minimax using a similar algorithmic scheme to achieve the minimax rate. Our algorithmic rate also applies with sub-Gaussian noise. We illustrate the utility of this theory with examples including multivariate monotone functions, linear functionals over ellipsoids, and Lipschitz classes.
1 Introduction
This paper investigates minimax rates for nonparametric regression models under convex constraints. We build on the techniques of Neykov (2022), in which a lower bound is established using Fano’s inequality, and then provide a universal scheme that matches this lower bound. The minimax rate turns out to depend closely on the local geometry of the constraint set, in particular, in terms of its local packing numbers. We are able to handle cases where the function in $\mathcal{F}$ do not have a uniform sup-norm bound. [^1] While we can handle function spaces unbounded in sup norm constraints we are always assuming that the function space has a bounded $L_{2}(\mathbb{P}_{X})$ diameter. Our algorithmic upper bound also applies when the noise term is sub-Gaussian.
1.1 Problem Formulation and Notation
Let $\mathcal{F}\subset L_{2}(\mathbb{P}_{X})$ be a convex and complete class of real-valued functions on $\mathcal{X}\subseteq\mathbb{R}^{p}$. Suppose we observe $\{(X_{i},Y_{i})\}_{i=1}^{n}$ where for each $i\in[n]$, $X_{i}$ is independently drawn from some distribution $\mathbb{P}_{X}$ on $\mathcal{X}$ and $Y_{i}=\bar{f}(X_{i})+\xi_{i}\in\mathcal{Y}\subseteq\mathbb{R}$ where $\xi_{i}\sim\mathcal{N}(0,\sigma^{2})$ are i.i.d., $\sigma>0$ some fixed constant (which does not scale with the sample size), and $\bar{f}\in\mathcal{F}$ the true regression function. We introduce the notation $\vec{Y}=(Y_{1},\dots,Y_{n}),\vec{X}=(X_{1},\dots,X_{n})$. We also consider sub-Gaussian noise, where we say a random variable $Z$ is sub-Gaussian with mean $\mu$ and variance proxy $\tau^{2}$ if $\mathbb{E}[\exp(t(Z-\mu)]\leq\exp(t^{2}\tau^{2}/2)$ for all $t>0$ and write $Z\sim\mathrm{SG}(\mu,\tau^{2})$. In the sub-Gaussian setting, we will assume $\xi_{i}\sim\mathrm{SG}(0,\sigma^{2})$ are drawn i.i.d.
For any $f,g\in\mathcal{F}$, define the $L_{2}(\mathbb{P}_{X})$-distance between $f$ and $g$ by
$\|f-g\|_{L_{2}(\mathbb{P}_{X})}=\sqrt{\int|f(x)-g(x)|^{2}\mathrm{d}\mathbb{P}_{X}(x)}.$
The $L_{p}(\mathbb{P}_{X})$-norm can be defined analagously for $p\geq 2$. We define $d:=\mathrm{diam}(\mathcal{F}):=\sup\{\|f-g\|_{L_{2}(\mathbb{P}_{X})}:f,g\in\mathcal{F}\}$; we assume this quantity is finite throughout the paper. The empirical $L_{2}$-distance given input points $X_{1},\dots,X_{n}\sim\mathbb{P}_{X}$ is defined by
$\|f-g\|_{L_{2}(\mathbb{P}_{n})}=n^{-1/2}\sqrt{\sum_{i=1}^{n}|f(X_{i})-g(X_{i})|^{2}}.$
As a short-hand, let $g(\vec{X})$ be the vector with $g(X_{i})$ in the $i$th component, and similarly for $f(\vec{X})$ and $\overline{f}(\vec{X}).$ Then $\|f-g\|_{L_{2}(\mathbb{P}_{n})}=n^{-1/2}\|f(\vec{X})-g(\vec{X})\|_{2}$, using the Euclidean $L_{2}$ norm for vectors in $\mathbb{R}^{n}$. Let $B(f,r,\|\cdot\|_{L_{2}(\mathbb{P}_{X})})=B(f,r)=\{g\in\mathcal{F}:\|f-g\|_{L_{2}(\mathbb{P}_{X})}\leq r\}$. Unless specified, we will always use the $L_{2}(\mathbb{P}_{X})$-norm for balls and abbreviate the notation as $B(f,r)$.
Our goal is to quantify the minimax rate
$\inf_{\widehat{f}}\sup_{\bar{f}\in\mathcal{F}}\mathbb{E}_{X,Y}\|\widehat{f}(\vec{X},\vec{Y})-\bar{f}\|_{L_{2}(\mathbb{P}_{X})}^{2}$
and achieve it with a universal scheme depending only on the local geometry of $\mathcal{F}$. Here the infimum is over all measurable functions of the data $\widehat{f}\colon\mathcal{X}^{n}\times\mathcal{Y}^{n}\to\mathcal{F}$.
An essential tool will be the local metric entropy of $\mathcal{F}$, which we now define. Observe that this quantity may be infinite if $\mathcal{F}$ is not totally bounded, but our theorems will still hold (see Remark 2.13).
Definition 1.1 (Local Entropy).
Let $M(r,\mathcal{F})$ be the largest cardinality of a $r$-packing set in the $L_{2}(\mathbb{P}_{X})$ norm of $\mathcal{F}$ (see Definition 5.4 Wainwright, 2019, e.g., for a definition of a packing set). Define for a constant $c>0$ the following:
$M_{\mathcal{F}}^{\operatorname{loc}}(\varepsilon,c)=\sup_{f\in\mathcal{F}}M(\varepsilon/c,B(f,\varepsilon)\cap\mathcal{F}).$
Then $\log M^{\operatorname{loc}}_{\mathcal{F}}(\varepsilon,c)$ is the local entropy of $\mathcal{F}$.
1.2 Related Works
The literature on nonparametric regression is vast (e.g., Tsybakov, 2009), so we do not give an exhaustive summary. We do however highlight those works closest in spirit to our results.
In Yang and Barron (1999), the authors study minimax rates in terms of global entropies for density estimation, but apply their results to nonparametric regression. They obtain a similar result in their Theorem 6, where the minimax rate is $\varepsilon^{2}$ for $\varepsilon$ satisfying $\log M(\varepsilon,\mathcal{F})\asymp n\varepsilon^{2}$. However they require $\sup_{f\in\mathcal{F}}\|f\|_{\infty}<L$ for some $L>0$, while our results relax this assumption. As a consequence, we are able to derive minimax rates on linear functionals with unbounded sup norms. In addition, yang1999information impose an assumption on the parameter space which essentially forces the local and global metric entropies to coincide. We do not require this condition in our work.
mendelson_2017_local_versus_global considers prediction problems for convex function classes, including ones with unbounded sup-norm. The minimax rate we consider can be viewed as a special case of this prediction setting. However, Mendelson bounds the error with high probability rather than in expectation. Moreover, mendelson_2017_local_versus_global requires an assumption that $\mathcal{F}$ is $L$-sub-Gaussian. This definition has two components. First, for each $p\geq 2$ and $f,g\in\mathcal{F}\cup\{0\}$, there must exist $L>0$ such that $\|f-g\|_{L_{p}(\mathbb{P}_{X})}\leq L\sqrt{p}\|f-g\|_{L_{2}(\mathbb{P}_{X})}$. Second, the centered canonical Gaussian process induced by $\mathcal{F}$ must be bounded. That is, to each $f\in\mathcal{F}$, we associate a Gaussian random variable $G_{f}$ with $\mathbb{E}[G_{f}]=0$ and $\mathbb{E}[G_{f}G_{g}]=\mathbb{E}[fg]$ for all $f,g\in\mathcal{F}$, and the assumption is that $\sup_{f}G_{f}$ is a bounded random variable almost surely. Our approach only uses the first assumption (and only for the unbounded sup-norm case) along with the diameter being bounded in the $L_{2}(\mathbb{P}_{X})$ norm. This assumption about the diameter may be easier to verify in practice compared to the boundedness of the Gaussian process. Another distinction with our approach is that mendelson_2017_local_versus_global does not match the lower and upper bounds on the error rate.
Other works include agglomeration_yang_2004, which examines taking data-dependent convex combinations of $n^{\tau}$ for $\tau>0$ many candidate regression procedures to obtain a rate of order of the minimax rate plus a penalty term. agglomeration_yang_2004 requires the function to be bounded in sup-norm, but does relax assumptions on the distribution of the error term. yang_2014_high_dimensional_nonparam studies the nonparametric regression setting under assumptions of sparsity or additivity of the function class. They also require assumptions that the distribution on the input space $\mathcal{X}=[0,1]^{p}$ has a density lower bounded on a sub-hypercube (i.e., the data is not from a strictly lower dimensional subspace). The paper relies on global metric entropy conditions related to the minimax rate, as in yang1999information.
1.3 Organization
Our paper is organized as follows. In Section 2 we derive the minimax rate assuming that there is a uniform bound on the sup-norm of functions in $\mathcal{F}.$ Section 3 relaxes this assumption and only assumes the $L_{2}(\mathbb{P}_{X})$-diameter of $\mathcal{F}$ is finite. Section 4 refines the error-rate of the algorithm’s output to be adaptive to the true $\bar{f}$.
We then derive in Section 5 numerous existing and new results for particular examples of $\mathcal{F}$. One of the simplest examples is the set of linear functionals, i.e., when $K\subset\mathbb{R}^{p}$ is compact and convex, we consider functions of the form $X^{T}\beta$ and draw data $X_{1},\dots,X_{n}$ from an isotropic sub-Gaussian distribution. For $K$ being the $\ell_{1}$-ball and assuming the true $\beta^{\ast}$ is $s$-sparse, we use Sudakov minoration along with properties of the Gaussian width of tangent cones stated in chandrasekaran2012convex and amelunxen2014living to derive a minimax rate of order $\varepsilon\asymp\sqrt{s\log(d/s)/n}$, matching results from raskutti2011minimax. Following the ellipse example in neykov2022minimax, we again consider sub-Gaussian design and linear functionals over ellipsoids and recover a similar rate. We compare this to results from pathak2023noisy.
We then use existing results about the metric entropy from yang1999information and gao_Wellner_monotone and recover the well-known minimax rate for monotone functions. We similarly derive minimax rates for functions with smoothness assumptions, a well-studied setting (stone_1980).
2 Bounded Case
The bounded case assumes there is some $B_{\mathcal{F}}>0$ such that $\|f\|_{\infty}\leq B_{\mathcal{F}}$ for all $f\in\mathcal{F}$. In this case, we have $d\leq 2B_{\mathcal{F}}$ by the triangle inequality. We implicitly assume $\mathcal{F}$ is totally-bounded, but Remark 2.13 details how this is not a necessary assumption, including in the later unbounded and adaptive settings.
The structure of this section is as follows. First we derive a lower bound on the minimax rate using Fano’s inequality. In particular, we find that the minimax rate is at least $\varepsilon^{2}$ up to constants where $\varepsilon$ satisfies $\log M_{\mathcal{F}}^{\operatorname{loc}}(\varepsilon,c)>4(n\varepsilon^{2}/(2\sigma^{2})\vee\log 2)$.
To upper bound the minimax rate, we first construct an algorithm that outputs an estimator achieving a squared error rate of $\varepsilon^{2}$ when $\varepsilon$ satisfies a (discretized) condition of the form $n\varepsilon^{2}>2\log M_{\mathcal{F}}^{\operatorname{loc}}\left(\overline{C}\varepsilon,c\right)\vee\log 2$ for some constant $\overline{C}$.
To tie together these two bounds, we show that either the minimax rate is $d^{2}$, in which case our algorithm clearly outputs a function in $\mathcal{F}$ achieving this error rate, or it has an error rate ${\varepsilon^{\ast}}^{2}$ satisfying both the lower and upper bound conditions mentioned above. In fact, we establish the rate is ${\varepsilon^{\ast}}^{2}\wedge d^{2}$, where
$\varepsilon^{\ast}:=\sup\{\varepsilon:n\varepsilon^{2}\leq\log M_{\mathcal{F}}^{\operatorname{loc}}(\varepsilon,c)\}$ (2.1)
for some sufficiently large constant $c$.
2.1 Lower Bound
The lower bound follows from Fano’s Inequality:
Lemma 2.1 (Fano’s inequality for nonparametric Regression).
Let $f_{1},\ldots,f_{m}\in\mathcal{F}$ be a collection be $\varepsilon$-separated functions, i.e., satisfying $\|f_{i}-f_{j}\|_{L_{2}(\mathbb{P}_{X})}^{2}>\varepsilon$ for each $i\neq j$. Suppose $J$ is uniformly distributed over the index set $[m]$, and $(Y_{i}|X_{i},J=j)=f_{j}(X_{i})+\xi_{i}$ where $\xi_{i}\sim N(0,\sigma^{2})$ for each $i\in[n],j\in[m]$. Then
$\inf_{\widehat{f}}\sup_{\bar{f}\in\mathcal{F}}\mathbb{E}_{X,Y}\|\widehat{f}(\vec{X},\vec{Y})-\bar{f}\|_{L_{2}(\mathbb{P}_{X})}^{2}\geq\frac{\varepsilon^{2}}{4}\bigg{(}1-\frac{I(\vec{X},\vec{Y};J)+\log 2}{\log m}\bigg{)},$
where $I(\vec{X},\vec{Y};J)$ is the mutual information between our data and $J$.
The next lemma gives an upper bound on the the mutual information term that appears in Fano’s Inequality.
Lemma 2.2 (Bounding the Mutual Information in Nonparametric Setting).
Under the same set-up as in Lemma 2.1, we have $I(\vec{X},\vec{Y};J)\leq\frac{n}{2\sigma^{2}}\max_{j}\|f_{j}-h\|_{L_{2}(\mathbb{P}_{X})}^{2}$ for any $h\in\mathcal{F}$.
{toappendix}
Proof of Lemma 2.2.
We use (wainwright2019high, (15.52)), the decoupling property of KL-divergence, and the decomposition of the joint distribution of the $(X_{i},Y_{i})$. Let $\mathbb{P}_{Y_{i}|X_{i}}^{f_{j}}$ be the distribution of $(Y_{i}|X_{i},J=j)$, which is a normal distribution with mean $f_{j}(X_{i})$ and variance $\sigma^{2}$. Denote by $\mathbb{P}_{(X_{i},Y_{i})}^{f_{j}}$ the distribution of $(Y_{i},X_{i}|J=j)$ and $\mathbb{P}_{X_{i}}$ that of $X_{i}$, so that $\mathbb{P}_{(X_{i},Y_{i})}^{f_{j}}=\mathbb{P}_{X_{i}}\times\mathbb{P}_{Y_{i}|X_{i}}^{f_{j}}$. Note that the $X_{i}$ does not depend on $f_{j}$. Similarly, for any function $h\in\mathcal{F}$, if we set $Y_{i}=h(X_{i})+\xi_{i}$ instead, denote the analogous distributions by $\mathbb{P}_{(X_{i},Y_{i})}^{h}=\mathbb{P}_{X_{i}}\times\mathbb{P}_{Y_{i}|X_{i}}^{h}$. Then we obtain
$I(\vec{X},\vec{Y};J)$ $\leq\frac{1}{m}\sum_{j}D_{KL}\left(\prod_{i}\mathbb{P}_{(X_{i},Y_{i})}^{f_{j}}\parallel\prod_{i}\mathbb{P}_{(X_{i},Y_{i})}^{h}\right)$
$=\frac{1}{m}\sum_{j}\sum_{i}D_{KL}\left(\mathbb{P}_{(X_{i},Y_{i})}^{f_{j}}\parallel\mathbb{P}_{(X_{i},Y_{i})}^{h}\right)$
$=\frac{1}{m}\sum_{j}\sum_{i}\mathbb{E}_{f_{j}}\left[\log\left(\mathbb{P}_{(X_{i},Y_{i})}^{f_{j}}/\mathbb{P}_{(X_{i},Y_{i})}^{h}\right)\right]$
$=\frac{1}{m}\sum_{j}\sum_{i}\mathbb{E}_{f_{j}}\left[\log\left(\frac{\mathbb{P}_{X_{i}}\mathbb{P}_{(Y_{i}|X_{i})}^{f_{j}}}{\mathbb{P}_{X_{i}}\mathbb{P}_{(Y_{i}|X_{i})}^{h}}\right)\right]$
$=\frac{1}{m}\sum_{j}\sum_{i}\mathbb{E}_{X_{i}}\mathbb{E}_{Y_{i}|(X_{i},f_{j})}\left[\log\left(\mathbb{P}_{(Y_{i}|X_{i})}^{f_{j}}/\mathbb{P}_{(Y_{i}|X_{i})}^{h}\right)\right]$
$=\frac{1}{m}\sum_{j}\sum_{i}\mathbb{E}_{X_{i}}\left[D_{KL}\left(N(f_{j}(X_{i}),\sigma^{2})\parallel N(h(X_{i}),\sigma^{2})\right)\right]$
$=\frac{1}{m}\sum_{j}\sum_{i}\mathbb{E}_{X_{i}}\left[\frac{|f_{j}(X_{i})-h(X_{i})|^{2}}{2\sigma^{2}}\right]$
$=\frac{1}{2\sigma^{2}m}\sum_{j}\sum_{i}\|f_{j}-h\|_{L_{2}(\mathbb{P}_{X})}^{2}$
$=\frac{n}{2\sigma^{2}m}\sum_{j}\|f_{j}-h\|_{L_{2}(\mathbb{P}_{X})}^{2}$
$\leq\frac{n}{2\sigma^{2}}\max_{j}\|f_{j}-h\|_{L_{2}(\mathbb{P}_{X})}^{2}.$
To complete the lower bound, we return to the local metric entropy defined earlier.
Lemma 2.3.
We have
$\inf_{\widehat{f}}\sup_{\bar{f}\in\mathcal{F}}\mathbb{E}_{X,Y}\|\widehat{f}(\vec{X},\vec{Y})-\bar{f}\|_{L_{2}(\mathbb{P}_{X})}^{2}\geq\frac{\varepsilon^{2}}{8c^{2}},$
for any $\varepsilon$ satisfying $\log M_{\mathcal{F}}^{\operatorname{loc}}(\varepsilon,c)>4\left(\frac{n\varepsilon^{2}}{2\sigma^{2}}\vee\log 2\right)$, where $c$ is the constant from the Definition 1.1 which is fixed to some large enough value.
{toappendix}
Proof of Lemma 2.3.
Pick a $f\in\mathcal{F}$ that achieves $M:=M_{\mathcal{F}}^{\operatorname{loc}}(\varepsilon,c):=\sup_{f\in\mathcal{F}}M(\varepsilon/c,B(f,\varepsilon)\cap\mathcal{F})$ (otherwise, repeat the following argument by picking a limiting sequence of functions). Let $f_{1},\dots,f_{M}$ be a maximal $\varepsilon/c$-packing set for $B(f,\varepsilon)\cap\mathcal{F}$, so that $\|f_{i}-f_{j}\|_{L_{2}(\mathbb{P}_{X})}>\varepsilon/c$ for all $i\neq j$. Because $f_{1},\dots,f_{M}\in B(f,\varepsilon)\cap\mathcal{F}$, it follows for each $j\in[M]$ that $\|f_{j}-f\|_{L_{2}(\mathbb{P}_{X})}\leq\varepsilon$. Letting $J$ be uniformly distributed on $[M]$, we have from Lemma 2.2 that
$I(\vec{X},\vec{Y};J)$ $\leq\frac{n}{2\sigma^{2}}\max_{j}\|f_{j}-f\|_{L_{2}(\mathbb{P}_{X})}^{2}\leq\frac{n\varepsilon^{2}}{2\sigma^{2}}.$
Applying Lemma 2.1, we obtain
$\inf_{\widehat{f}}\sup_{\bar{f}\in\mathcal{F}}\mathbb{E}_{X,Y}\|\widehat{f}(\vec{X},\vec{Y})-\bar{f}\|_{L_{2}(\mathbb{P}_{X})}^{2}$ $\geq\frac{\varepsilon^{2}}{4c^{2}}\bigg{(}1-\frac{\frac{n\varepsilon^{2}}{2\sigma^{2}}+\log 2}{\log M}\bigg{)}.$
Choose $\varepsilon$ such that $\log M$ is bigger than the middle term in
$\log M\geq 4\left(\frac{n\varepsilon^{2}}{2\sigma^{2}}\vee\log 2\right)\geq 2\left(\frac{n\varepsilon^{2}}{2\sigma^{2}}+\log 2\right),$
and this proves the minimax risk is lower-bounded by $\frac{\varepsilon^{2}}{8c^{2}}$. ∎
2.2 Upper Bound
We now propose Algorithm 1, which produces an $f^{\ast}$ achieving the minimax rate up to constants, using a procedure similar to that in neykov2022minimax. We begin by providing an informal description. We will take $c=2(C+1)$ where $C>3$ is some universal constant. In the following section, all distances are with respect to the $L_{2}(\mathbb{P}_{X})$-norm.
Pick an arbitrary $f^{\ast}\in\mathcal{F}$. Recalling that $d$ is the diameter of $\mathcal{F}$, form a maximal $\frac{d}{c}$-packing set $M_{1}$ of the set $B(f^{\ast},d)\cap\mathcal{F}$. We then update $f^{\ast}$ by reassigning it to be the least squares estimator over the $|M_{1}|$ functions in $M_{1}$:
$f^{\ast}=\mathop{\mathrm{argmin}}_{f\in M_{1}}\sum_{i=1}^{n}(Y_{i}-f(X_{i}))^{2}.$
The process now repeats but with a smaller ball and packing set (divide $d/c$ or $d$, respectively, by the next power of 2). That is, we now form a maximal $\frac{d}{2c}$-packing set $M_{2}$ of the set $B(f^{\ast},d/2)\cap\mathcal{F}$ and reassign
$f^{\ast}=\mathop{\mathrm{argmin}}_{f\in M_{2}}\sum_{i=1}^{n}(Y_{i}-f(X_{i}))^{2}.$
Then iterate the process for as many steps as desired, e.g., until some stopping condition is attained. Note that by construction of the packing sets, the updates form a Cauchy-sequence and the outputs will converge. To see this, denote by $\Upsilon_{j}$ the $j$th update of $f^{\ast}$, where $\Upsilon_{1}$ is our starting choice. Then for $j<k$, we have
$\|\Upsilon_{j}-\Upsilon_{k}\|_{L_{2}(\mathbb{P}_{X})}$ $\leq\sum_{i=j}^{k-1}\|\Upsilon_{i+1}-\Upsilon_{i}\|_{L_{2}(\mathbb{P}_{X})}\leq\sum_{i=j}^{k-1}\frac{d}{2^{i-1}}\leq\frac{d}{2^{j-2}}.$ (2.2)
Thus for any $\varepsilon>0$, for sufficiently large $N$, if $j,k\geq N$, we have $\|\Upsilon_{j}-\Upsilon_{k}\|_{L_{2}(\mathbb{P}_{X})}\leq\varepsilon$. Recall moreover that $\mathcal{F}$ is complete. The packing sets may become infinite if $\mathcal{F}$ is not totally bounded. But the algorithm’s output remains a Cauchy sequence.
The construction of the packing sets is done in a data-independent manner by forming an infinite tree of possible updates of $f^{\ast}$. After forming the first packing set, the sample space $\mathcal{X}^{n}\times\mathcal{Y}^{n}$ is partitioned into $|M_{1}|$ regions corresponding to which of the $|M_{1}|$ functions is the least squares estimator. For each possible choice of updated $f^{\ast}$, a new packing set is formed, and again the sample space is partitioned. Performing the algorithm on a particular data realization corresponds to traversing a single (infinite) path down the tree.
This procedure may encounter ties in the minimization step, so we must modify our algorithm to handle such cases and in doing so, preserve measurability of the estimator it outputs. Our goal is to induce a lexicographic ordering on our functions and then choose the smallest function in the ordering to break ties. The details are given in the appendix (Remark 2.4).
{toappendix}
Remark 2.4 (Breaking Ties in Algorithm).
By Theorem A.1 in gyorfi_2022, the set of continuous functions is dense in $\mathcal{F}$. Hence we can construct all of our packing sets using continuous functions only. That is, if $\mathcal{F}_{c}$ denote the set of all continuous functions in $\mathcal{F}$, then we construct $M_{k}$ as the maximal packing of $B(f^{*},d/2^{k-1})\cap\mathcal{F}_{c}$. This $M_{k}$ clearly has cardinality that is smaller than or equal to the $M_{k}$ in our previous notation since $\mathcal{F}_{c}\subseteq\mathcal{F}$. In addition since the set of all continuous functions is dense in $\mathcal{F}$, we have that for any $g\in\mathcal{F}$ there exists an $h\in\mathcal{F}_{c}$ such that $\|g-h\|\leq\varepsilon$ for any $\varepsilon>0$. Pick the point from $f\in M_{k}$ which is $d/2^{k-1}$ close to $h$ and conclude that $\|f-g\|\leq\varepsilon+d/2^{k-1}$. Let $\varepsilon$ go to zero, and we conclude that any point in $\mathcal{F}$ can be approximated by a point in $\mathcal{F}_{c}$ within distance $d/2^{k-1}$. Now suppose the domain of the functions is $\mathbb{R}^{p}$. Take the set $\mathbb{Q}^{p}$ which is dense in $\mathbb{R}^{p}$. This is a well ordered set since there is a bijection between $\mathbb{Q}^{p}$ and $\mathbb{N}$. Hence we can look at the functions from $M_{k}$ and their restrictions to $\mathbb{Q}^{p}$ and order them lexicographically. There will not be any ties, since the functions are separated and continuous (hence the functions cannot be the same on $\mathbb{Q}^{p}$ otherwise they are the same on $\mathbb{R}^{p}$ which is a contradiction) so we can take the smallest function in the lexicographic ordering.
Input: A point $f^{\ast}\in\mathcal{F}$, norm $L_{2}(\mathbb{P}_{X})$, step count $T$
1 $k\leftarrow 1$;
$\Upsilon\leftarrow[f^{\ast}\in\mathcal{F}]$ ;
  /* This array is needed solely in the proof and is not used by the estimator */
2 while $k\leq T-1$ do
       Take a $\frac{d}{2^{k}(C+1)}$ maximal packing set $M_{k}$ of the set $B\big{(}f^{\ast},\frac{d}{2^{k-1}}\big{)}\cap\mathcal{F}$ ;
        /* The packing sets should be constructed prior to seeing the data */
       $f^{\ast}\leftarrow\mathop{\mathrm{argmin}}_{f\in M_{k}}\sum_{i=1}^{n}(Y_{i}-f(X_{i}))^{2}$ ;
        /* Break ties by picking the point with the least lexicographic ordering. */
3       $\Upsilon$.append$(f^{\ast})$;
4       $k\leftarrow k+1$;
5
return $f^{\ast}$
Algorithm 1 Upper Bound Algorithm for Nonparametric Regression
To prove this algorithm has the desired properties, we first relate the $L_{2}(\mathbb{P}_{X})$ norm to the $L_{2}(\mathbb{P}_{n})$ norm. In particular, Lemma 2.5 shows that if $f,g\in\mathcal{F}$ are far apart but $f$ is close to $\bar{f}$ in $L_{2}(\mathbb{P}_{X})$-distances, then the same is true for $f,g,\bar{f}$ with high probability in the $L_{2}(\mathbb{P}_{n})$-distance. The proof, given in the appendix, relies on Bernstein’s concentration inequality. This result also applies when the noise is sub-Gaussian.
Lemma 2.5.
Suppose $f,g\in\mathcal{F}$ satisfy $n\|f-g\|_{L_{2}(\mathbb{P}_{X})}^{2}\geq C^{2}\delta^{2}$ for some sufficiently large $C>0$. Suppose further that $n\|f-\bar{f}\|_{L_{2}(\mathbb{P}_{X})}^{2}<\delta^{2}$. Then $n\|f-\bar{f}\|_{L_{2}(\mathbb{P}_{n})}^{2}\leq 2\delta^{2}$ and $n\|f-g\|_{L_{2}(\mathbb{P}_{n})}^{2}\geq(C^{2}-1)\delta^{2}$ both simultaneously hold with probability at least
$1-\exp\left(-\frac{3\delta^{2}}{32B_{\mathcal{F}}^{2}}\right)-\exp\left(-\frac{3\delta^{2}}{8B_{\mathcal{F}}^{2}\cdot(3C^{2}+1)}\right).$
{toappendix}
Proof of Lemma 2.5.
We recall Bernstein’s inequality (bennetbernstein1962, see pages 33-34). Namely, given i.i.d. random variables $Z_{1},\dots,Z_{n}$ with $\mathbb{E}[Z_{1}]=0$, $|Z_{1}|\leq K$ almost surely, and any $t>0$, we have
$\mathbb{P}\left(\sum_{i}Z_{i}>t\right)$ $\leq\exp\left(-\frac{t^{2}/2}{n\mathbb{E}[Z_{1}^{2}]+Kt/3}\right).$
If we replace the $Z_{i}$ with $-Z_{i}$, a similar one-sided tail bound is derived, with $t>0$ still:
$\mathbb{P}\left(\sum_{i}Z_{i}<-t\right)$ $\leq\exp\left(-\frac{t^{2}/2}{n\mathbb{E}[Z_{1}^{2}]+Kt/3}\right).$
We set $Z_{i}=|f(X_{i})-g(X_{i})|^{2}-\|f-g\|_{L_{2}(\mathbb{P}_{X})}^{2}$ for each $i\in[n]$, which are mean 0 and i.i.d. random variables with $|Z_{1}|\leq K$ for $K=8B_{\mathcal{F}}^{2}$, so that $\sum_{i=1}^{n}Z_{i}=n\|f-g\|_{L_{2}(\mathbb{P}_{n})}^{2}-n\|f-g\|_{L_{2}(\mathbb{P}_{X})}^{2}.$ Moreover,
$\mathbb{E}[Z_{1}^{2}]$ $\leq\mathbb{E}|f(X_{i})-g(X_{i})|^{4}\leq 4B_{\mathcal{F}}^{2}\mathbb{E}|f(X_{i})-g(X_{i})|^{2}=K\|f-g\|_{L_{2}(\mathbb{P}_{X})}^{2}.$
We set $t=C^{-2}\cdot n\|f-g\|_{L_{2}(\mathbb{P}_{X})}^{2}$. The first and last inequality use $n\|f-g\|_{L_{2}(\mathbb{P}_{X})}^{2}\geq C^{2}\delta^{2}$, and the second inequality is due to Bernstein.
$\mathbb{P}_{\bar{f}}\left(n\|f-g\|_{L_{2}(\mathbb{P}_{n})}^{2}<(C^{2}-1)\delta^{2}\right)$ $=\mathbb{P}_{\bar{f}}\left(n\|f-g\|_{L_{2}(\mathbb{P}_{n})}^{2}<(1-C^{-2})\cdot C^{2}\delta^{2}\right)$
$\leq\mathbb{P}_{\bar{f}}\left(n\|f-g\|_{L_{2}(\mathbb{P}_{n})}^{2}<(1-C^{-2})\cdot n\|f-g\|_{L_{2}(\mathbb{P}_{X})}^{2}\right)$
$=\mathbb{P}_{\bar{f}}\left(n\|f-g\|_{L_{2}(\mathbb{P}_{n})}^{2}-n\|f-g\|_{L_{2}(\mathbb{P}_{X})}^{2}<\right.$
$\quad\quad\quad\quad\quad\left.-C^{-2}\cdot n\|f-g\|_{L_{2}(\mathbb{P}_{X})}^{2}\right)$
$=\mathbb{P}_{\bar{f}}\left(\sum_{i=1}^{n}Z_{i}<-t\right)$
$\leq\exp\left(-\frac{t^{2}/2}{n\mathbb{E}[Z_{1}^{2}]+Kt/3}\right)$
$=\exp\left(-\frac{t^{2}/2}{Kn\|f-g\|_{L_{2}(\mathbb{P}_{X})}^{2}+Kt/3}\right)$
$=\exp\left(-\frac{3t^{2}}{6Kn\|f-g\|_{L_{2}(\mathbb{P}_{X})}^{2}+2Kt}\right)$
$=\exp\left(-\frac{3t^{2}}{6KtC^{2}+2Kt}\right)$
$=\exp\left(-\frac{3t}{2K(3C^{2}+1)}\right)$
$=\exp\left(-\frac{3n\|f-g\|_{L_{2}(\mathbb{P}_{X})}^{2}\cdot C^{-2}}{16B_{\mathcal{F}}^{2}(3C^{2}+1)}\right)$
$\leq\exp\left(-\frac{3C^{2}\delta^{2}\cdot C^{-2}}{16B_{\mathcal{F}}^{2}(3C^{2}+1)}\right)$
$=\exp\left(-\frac{3\delta^{2}}{16B_{\mathcal{F}}^{2}(3C^{2}+1)}\right).$
Thus, we have shown that $n\|f-g\|_{L_{2}(\mathbb{P}_{n})}^{2}\geq(C^{2}-1)\delta^{2}$ with probability at least
$1-\exp\left(-\frac{3\delta^{2}}{16B_{\mathcal{F}}^{2}\cdot(3C^{2}+1)}\right).$
For the other bound, we now set $Z_{i}=|f(X_{i})-\bar{f}(X_{i})|^{2}-\|f-\bar{f}\|_{L_{2}(\mathbb{P}_{X})}^{2}$ for each $i\in[n]$, which are again mean 0 and i.i.d. random variables with $|Z_{1}|\leq K$ for $K=8B_{\mathcal{F}}^{2}$, so that $\sum_{i=1}^{n}Z_{i}=n\|f-\bar{f}\|_{L_{2}(\mathbb{P}_{n})}^{2}-n\|f-\bar{f}\|_{L_{2}(\mathbb{P}_{X})}^{2}.$ Then similarly,
$\mathbb{E}[Z_{1}^{2}]$ $\leq\mathbb{E}|f(X_{i})-\bar{f}(X_{i})|^{4}\leq 4B_{\mathcal{F}}^{2}\mathbb{E}|f(X_{i})-\bar{f}(X_{i})|^{2}=K\|f-\bar{f}\|_{L_{2}(\mathbb{P}_{X})}^{2}.$
This implies $n\mathbb{E}[Z_{1}^{2}]\leq K\delta^{2}$. Hence using $t=\delta^{2}$,
$\mathbb{P}_{\bar{f}}\left(n\|f-\bar{f}\|_{L_{2}(\mathbb{P}_{n})}^{2}>2\delta^{2}\right)$ $\leq\mathbb{P}_{\bar{f}}\left(n\|f-\bar{f}\|_{L_{2}(\mathbb{P}_{n})}^{2}>n\|f-\bar{f}\|_{L_{2}(\mathbb{P}_{X})}^{2}+\delta^{2}\right)$
$\leq\exp\left(-\frac{\delta^{4}/2}{n\mathbb{E}[Z_{1}^{2}]+K\delta^{2}/3}\right)$
$\leq\exp\left(-\frac{\delta^{4}/2}{K\delta^{2}+K\delta^{2}/3}\right)$
$=\exp\left(-\frac{3\delta^{2}}{64B_{\mathcal{F}}^{2}}\right).$
Thus, we have shown that $n\|f-\bar{f}\|_{L_{2}(\mathbb{P}_{n})}^{2}\leq 2\delta^{2}$ with probability at least
$1-\exp\left(-\frac{3\delta^{2}}{64B_{\mathcal{F}}^{2}}\right).$
Combining our results with a union bound, with probability at least
$1-\exp\left(-\frac{3\delta^{2}}{64B_{\mathcal{F}}^{2}}\right)-\exp\left(-\frac{3\delta^{2}}{16B_{\mathcal{F}}^{2}\cdot(3C^{2}+1)}\right),$
we have both $n\|f-\bar{f}\|_{L_{2}(\mathbb{P}_{n})}^{2}\leq 2\delta^{2}$ and $n\|f-g\|_{L_{2}(\mathbb{P}_{n})}^{2}\geq(C^{2}-1)\delta^{2}$.
The next lemma considers hypothesis testing whether the true function is $f$ or $g$ when $f,g\in\mathcal{F}$ are at least $L_{2}(\mathbb{P}_{X})$-distance $\delta/\sqrt{n}$ apart. When $f$ is $\delta/\sqrt{n}$ close to the truth, the probability our test claims $g$ is the right choice is low, and similarly, when $g$ is close to the truth, the probability our test claims $f$ is the right choice is low.
Lemma 2.6.
Suppose we are testing $H_{0}:\bar{f}=f$ vs $H_{1}:\bar{f}=g$ where $n\|f-g\|_{L_{2}(\mathbb{P}_{X})}^{2}\geq C^{2}\delta^{2}$ for some $C>3$. Then the test $\psi(\vec{Y})=\mathbbm{1}\left(\|\vec{Y}-f(\vec{X})\|_{2}^{2}\geq\|\vec{Y}-g(\vec{X})\|_{2}^{2}\right)$ satisfies
$\sup_{\bar{f}:n\|f-\bar{f}\|_{L_{2}(\mathbb{P}_{X})}^{2}<\delta^{2}}\mathbb{P}_{\bar{f}}(\psi=1)$ $\vee\sup_{\bar{f}:n\|g-\bar{f}\|_{L_{2}(\mathbb{P}_{X})}^{2}<\delta^{2}}\mathbb{P}_{\bar{f}}(\psi=0)\leq 3\exp\left(-L(C,\sigma,B_{\mathcal{F}})\cdot\delta^{2}\right),$
where
$L(C,\sigma,B_{\mathcal{F}})=\min\left\{\frac{\left(\sqrt{C^{2}-1}-2\sqrt{2}\right)^{2}}{8\sigma^{2}},\frac{3}{64B_{\mathcal{F}}^{2}},\frac{3}{16B_{\mathcal{F}}^{2}\cdot(3C^{2}+1)}\right\}.$ (2.3)
{toappendix}
Proof of Lemma 2.6.
Let $\eta(\vec{X})=\bar{f}(\vec{X})-f(\vec{X})$. Then we write
$\|\vec{Y}-f(\vec{X})\|_{2}^{2}-\|\vec{Y}-g(\vec{X})\|_{2}^{2}$ $=\sum_{i=1}^{n}(Y_{i}-f(X_{i}))^{2}-\sum_{i=1}^{n}(Y_{i}-g(X_{i}))^{2}$
$=2\sum_{i=1}^{n}Y_{i}(g(X_{i})-f(X_{i}))+\sum_{i=1}^{n}f(X_{i})^{2}-\sum_{i=1}^{n}g(X_{i})^{2}$
$=2\sum_{i=1}^{n}(\bar{f}(X_{i})+\xi_{i})(g(x_{i})-f(x_{i}))+\|f(\vec{X})\|_{2}^{2}-\|g(\vec{X})\|_{2}^{2}$
$=2(\bar{f}(\vec{X})+\xi)^{T}(g(\vec{X})-f(\vec{X}))+\|f(\vec{X})\|_{2}^{2}-\|g(\vec{X})\|_{2}^{2}$
$=2(f(\vec{X})+\eta(\vec{X})+\xi)^{T}(g(\vec{X})-f(\vec{X}))+\|f(\vec{X})\|_{2}^{2}-\|g(\vec{X})\|_{2}^{2}$
$=2f(\vec{X})^{T}(g(\vec{X})-f(\vec{X}))+2\eta^{T}(g(\vec{X})-f(\vec{X}))$
$\quad\quad\quad 2\xi(\vec{X})^{T}(g(\vec{X})-f(\vec{X}))+\|f(\vec{X})\|_{2}^{2}-\|g(\vec{X})\|_{2}^{2}$
$=-\|f(\vec{X})-g(\vec{X})\|_{2}^{2}+2\eta(\vec{X})^{T}(g(\vec{X})-f(\vec{X}))$
$\quad\quad\quad+2\xi^{T}(g(\vec{X})-f(\vec{X})).$
Observe that conditional on $\vec{X}$, by applying the Cauchy-Schwarz inequality, this quantity is normal with mean at most
$-\|f(\vec{X})-g(\vec{X})\|_{2}^{2}+2\|\eta(\vec{X})\|_{2}\|f(\vec{X})-g(\vec{X})\|_{2}$
and variance $4n\sigma^{2}\|f-g\|_{L_{2}(\mathbb{P}_{n})}^{2}$.
We now derive estimates on $\|\eta(\vec{X})\|_{2}$ and $\|f-g\|_{L_{2}(\mathbb{P}_{n})}^{2}$ that hold with high probability. Suppose $n\|f-\bar{f}\|_{L_{2}(\mathbb{P}_{X})}^{2}<\delta^{2}$. From Lemma 2.5, this condition along with our assumption $n\|f-g\|_{L_{2}(\mathbb{P}_{X})}^{2}\geq C^{2}\delta^{2}$ implies that with probability at least $1-\exp\left(-\frac{3\delta^{2}}{64B_{\mathcal{F}}^{2}}\right)-\exp\left(-\frac{3\delta^{2}}{16B_{\mathcal{F}}^{2}\cdot(3C^{2}+1)}\right)$, we have
$n\|f-g\|_{L_{2}(\mathbb{P}_{n})}^{2}\geq(C^{2}-1)\delta^{2},\quad n\|f-\bar{f}\|_{L_{2}(\mathbb{P}_{n})}^{2}\leq 2\delta^{2}.$ (2.4)
Let $E$ denote the event that both inequalities of (2.4) hold, so
$\mathbb{P}_{\bar{f}}(E^{c})\leq\exp\left(-\frac{3\delta^{2}}{64B_{\mathcal{F}}^{2}}\right)+\exp\left(-\frac{3\delta^{2}}{16B_{\mathcal{F}}^{2}\cdot(3C^{2}+1)}\right).$
The first inequality in (2.4) implies
$\|f(\vec{X})-g(\vec{X})\|_{2}=\sqrt{n}\|f-g\|_{L_{2}(\mathbb{P}_{n})}\geq\delta\sqrt{C^{2}-1}$
in which case
$\delta\leq\frac{\|f(\vec{X})-g(\vec{X})\|_{2}}{\sqrt{C^{2}-1}}.$
The other inequality simply states that $\|\eta(\vec{X})\|_{2}\leq\delta\sqrt{2}$. Thus, the inequalities of (2.4) imply
$\|\eta(\vec{X})\|_{2}\leq\delta\sqrt{2}\leq\frac{\sqrt{2}}{\sqrt{C^{2}-1}}\cdot\|f(\vec{X})-g(\vec{X})\|_{2}$
Thus, conditional on $\vec{X}$, the event $E$ implies $\|\vec{Y}-f(\vec{X})\|_{2}^{2}-\|\vec{Y}-g(\vec{X})\|_{2}^{2}$ is normal with mean bounded from above by
$-\|f(\vec{X})-g(\vec{X})\|_{2}^{2}+2\|\eta(\vec{X})\|_{2}\|f(\vec{X})-g(\vec{X})\|_{2}$ $\leq\left(-1+\frac{2\sqrt{2}}{\sqrt{C^{2}-1}}\right)\cdot\|f(\vec{X})-g(\vec{X})\|_{2}^{2}$
$=\underbrace{\left(-1+\frac{2\sqrt{2}}{\sqrt{C^{2}-1}}\right)}_{<0}\cdot n\|f-g\|_{L_{2}(\mathbb{P}_{n})}^{2}$
and variance $4n\sigma^{2}\|f-g\|_{L_{2}(\mathbb{P}_{n})}^{2}$. Note that $\frac{2\sqrt{2}}{\sqrt{C^{2}-1}}<1$ because $C>3$.
Using a bound from van1996weak, we note that for any $\rho<0<\tau$,
$\mathbb{P}(N(\rho,\tau^{2})\geq 0)\leq\exp(-\rho^{2}/(2\tau^{2})).$
So conditional on $\vec{X}$ and the event $E$, $\psi=1$ holds with probability bounded by
$\exp\left(-\frac{\left(-1+\frac{2\sqrt{2}}{\sqrt{C^{2}-1}}\right)^{2}n^{2}\|f-g\|_{L_{2}(\mathbb{P}_{n})}^{4}}{8n\sigma^{2}\|f-g\|_{L_{2}(\mathbb{P}_{n})}^{2}}\right)$ $=\exp\left(-\frac{\left(-1+\frac{2\sqrt{2}}{\sqrt{C^{2}-1}}\right)^{2}n\|f-g\|_{L_{2}(\mathbb{P}_{n})}^{2}}{8\sigma^{2}}\right)$
$\leq\exp\left(-\frac{\left(-1+\frac{2\sqrt{2}}{\sqrt{C^{2}-1}}\right)^{2}(C^{2}-1)\delta^{2}}{8\sigma^{2}}\right)$
$=\exp\left(-\frac{\left(\sqrt{C^{2}-1}-2\sqrt{2}\right)^{2}\delta^{2}}{8\sigma^{2}}\right)$
Since $\mathbb{P}_{\bar{f}}(\psi=1|\vec{X},E)$ has an upper bound not depending on $X$, it follows (by taking an expectation) that
$\mathbb{P}_{\bar{f}}(\psi=1|E)\leq\exp\left(-\frac{\left(\sqrt{C^{2}-1}-2\sqrt{2}\right)^{2}\delta^{2}}{8\sigma^{2}}\right).$
Thus,
$\mathbb{P}_{\bar{f}}(\psi=1)$ $=\mathbb{P}_{\bar{f}}(\psi=1|E)\cdot\mathbb{P}_{\bar{f}}(E)+\mathbb{P}_{\bar{f}}(\psi=1|E^{c})\cdot\mathbb{P}_{\bar{f}}(E^{c})$
$\leq\mathbb{P}_{\bar{f}}(\psi=1|E)+\mathbb{P}_{\bar{f}}(E^{c})$
$\leq\exp\left(-\frac{\left(\sqrt{C^{2}-1}-2\sqrt{2}\right)^{2}\delta^{2}}{8\sigma^{2}}\right)+\exp\left(-\frac{3\delta^{2}}{64B_{\mathcal{F}}^{2}}\right)+\exp\left(-\frac{3\delta^{2}}{16B_{\mathcal{F}}^{2}\cdot(3C^{2}+1)}\right).$
If we instead assumed $n\|g-\bar{f}\|_{L_{2}(\mathbb{P}_{X})}^{2}<\delta^{2}$, a completely symmetric argument would result in a identical bound for $\mathbb{P}_{\bar{f}}(\psi=0)$. Thus, we conclude that
$\sup_{\bar{f}:n\|f-\bar{f}\|_{L_{2}(\mathbb{P}_{X})}^{2}<\delta^{2}}\mathbb{P}_{\bar{f}}(\psi=1)$ $\vee\sup_{\bar{f}:n\|g-\bar{f}\|_{L_{2}(\mathbb{P}_{X})}^{2}<\delta^{2}}\mathbb{P}_{\bar{f}}(\psi=0)\leq\ \exp\left(-\frac{\left(\sqrt{C^{2}-1}-2\sqrt{2}\right)^{2}\delta^{2}}{8\sigma^{2}}\right)$
$\quad\quad\quad+\exp\left(-\frac{3\delta^{2}}{64B_{\mathcal{F}}^{2}}\right)+\exp\left(-\frac{3\delta^{2}}{16B_{\mathcal{F}}^{2}\cdot(3C^{2}+1)}\right).$
We define
$L(C,\sigma,B_{\mathcal{F}})=\min\left\{\frac{\left(\sqrt{C^{2}-1}-2\sqrt{2}\right)^{2}}{8\sigma^{2}},\frac{3}{64B_{\mathcal{F}}^{2}},\frac{3}{16B_{\mathcal{F}}^{2}\cdot(3C^{2}+1)}\right\},$
to conclude that
$\sup_{\bar{f}:n\|f-\bar{f}\|_{L_{2}(\mathbb{P}_{X})}^{2}<\delta^{2}}\mathbb{P}_{\bar{f}}(\psi=1)$ $\vee\sup_{\bar{f}:n\|g-\bar{f}\|_{L_{2}(\mathbb{P}_{X})}^{2}<\delta^{2}}\mathbb{P}_{\bar{f}}(\psi=0)\leq 3\exp\left(-L(C,\sigma,B_{\mathcal{F}})\cdot\delta^{2}\right).$
We now build on the previous lemma by showing the function in a $\delta$-packing closest to our data will also be $\delta$-close (in the $L_{2}(\mathbb{P}_{X})$-distance) to the true function with high probability.
Lemma 2.7.
Let $f_{1},\dots,f_{M}$ be a $\delta$-packing of $\mathcal{F}^{\prime}\subseteq\mathcal{F}$ and suppose $\bar{f}\in\mathcal{F}^{\prime}$. Let $i^{\ast}=\mathop{\mathrm{argmin}}_{i\in[M]}\sum_{j=1}^{n}(Y_{j}-f_{i}(X_{j}))^{2}.$ Then
$\mathbb{P}_{\bar{f}}\left(\|f_{i^{\ast}}-\bar{f}\|_{L_{2}(\mathbb{P}_{X})}>\delta(C+1)\right)$ $\leq 3M\exp\left(-L(C,\sigma,B_{\mathcal{F}})\cdot n\delta^{2}\right),$
where $L(C,\sigma,B_{\mathcal{F}})$ is defined as in (2.3).
{toappendix}
Proof of Lemma 2.7.
For some $i\in[M]$, possibly distinct from $i^{\ast}$, we have $\|f_{i}-\bar{f}\|_{L_{2}(\mathbb{P}_{X})}\leq\delta$ which means $n\|f_{i}-\bar{f}\|_{L_{2}(\mathbb{P}_{X})}^{2}\leq n\delta^{2}$. By the triangle inequality, we observe
$\|f_{i^{\ast}}-\bar{f}\|_{L_{2}(\mathbb{P}_{X})}$ $\leq\|f_{i^{\ast}}-f_{i}\|_{L_{2}(\mathbb{P}_{X})}+\|f_{i}-\bar{f}\|_{L_{2}(\mathbb{P}_{X})}\leq\|f_{i^{\ast}}-f_{i}\|_{L_{2}(\mathbb{P}_{X})}+\delta.$
Note that $\|\vec{Y}-f_{i}(\vec{X})\|_{2}^{2}\geq\|\vec{Y}-f_{i^{\ast}}(\vec{X})\|_{2}^{2}$ since $f_{i^{\ast}}$ was defined to be the least squares estimator.
Putting these facts together,
$\mathbb{P}_{\bar{f}}\left(\|f_{i^{\ast}}-\bar{f}\|_{L_{2}(\mathbb{P}_{X})}>\delta(C+1)\right)$ $\leq\mathbb{P}_{\bar{f}}\left(\|f_{i^{\ast}}-f_{i}\|_{L_{2}(\mathbb{P}_{X})}>\delta\cdot C\right)$
$\leq\mathbb{P}_{\bar{f}}\left(\exists j:\|f_{i}-f_{j}\|_{L_{2}(\mathbb{P}_{X})}>\delta\cdot C\text{ and }\right.$
$\quad\quad\quad\quad\left.\|\vec{Y}-f_{i}(\vec{X})\|_{2}^{2}\geq\|\vec{Y}-f_{j}(\vec{X})\|_{2}^{2}\right)$
$=\mathbb{P}_{\bar{f}}\left(\exists j:n\|f_{i}-f_{j}\|_{L_{2}(\mathbb{P}_{X})}^{2}>C^{2}n\delta^{2}\text{ and }\right.$
$\quad\quad\quad\quad\left.\|\vec{Y}-f_{i}(\vec{X})\|_{2}^{2}\geq\|\vec{Y}-f_{j}(\vec{X})\|_{2}^{2}\right)$
$\leq\sum_{j}\mathbb{P}_{\bar{f}}\left(n\|f_{i}-f_{j}\|_{L_{2}(\mathbb{P}_{X})}^{2}>C^{2}n\delta^{2}\text{ and }\right.$
$\quad\quad\quad\quad\left.\|\vec{Y}-f_{i}(\vec{X})\|_{2}^{2}\geq\|\vec{Y}-f_{j}(\vec{X})\|_{2}^{2}\right)$
$\leq 3M\exp\left(-L(C,\sigma,B_{\mathcal{F}})\cdot n\delta^{2}\right).$ (2.5)
where we used Lemma 2.6 in (2.5) with $\sqrt{n}\delta$ in place of $\delta$.
Remark 2.8.
Lemma 2.6 and Lemma 2.7 can be generalized to our sub-Gaussian setting where each $\xi_{i}\sim\mathrm{SG}(0,\sigma^{2})$. We supply the details in the appendix.
{toappendix}
Proof of Remark 2.8.
In the proof of Lemma 2.6, we first showed that
$\|\vec{Y}-f(\vec{X})\|_{2}^{2}-\|\vec{Y}-g(\vec{X})\|_{2}^{2}$ $=-\|f(\vec{X})-g(\vec{X})\|_{2}^{2}+2\eta(\vec{X})^{T}(g(\vec{X})-f(\vec{X}))$
$\quad\quad\quad+2\xi^{T}(g(\vec{X})-f(\vec{X})).$
Conditional on $\vec{X}$, this quantity is sub-Gaussian with mean bounded by
$-\|f(\vec{X})-g(\vec{X})\|_{2}^{2}+2\|\eta(\vec{X})\|_{2}\|f(\vec{X})-g(\vec{X})\|_{2}.$
Moreover, the variance proxy can be shown to be $4n\sigma^{2}\|f-g\|_{L_{2}(\mathbb{P}_{n})}^{2}$ using the definition of a sub-Gaussian random variable. The proof of Lemma 2.6 then defines the event $E$ as when the inequalities in (2.4) hold, and $E$ satisfies
$\mathbb{P}_{\bar{f}}(E^{c})\leq\exp\left(-\frac{3\delta^{2}}{64B_{\mathcal{F}}^{2}}\right)+\exp\left(-\frac{3\delta^{2}}{16B_{\mathcal{F}}^{2}\cdot(3C^{2}+1)}\right).$
Noting that Lemma 2.5 still holds with sub-Gaussian noise, we obtain as in Lemma 2.6 that conditional on $\vec{X}$, the event $E$ implies $\|\vec{Y}-f(\vec{X})\|_{2}^{2}-\|\vec{Y}-g(\vec{X})\|_{2}^{2}$ is sub-Gaussian with mean bounded above by the negative quantity
$\left(-1+\frac{2\sqrt{2}}{\sqrt{C^{2}-1}}\right)\cdot n\|f-g\|_{L_{2}(\mathbb{P}_{n})}^{2}$
and variance proxy $4n\sigma^{2}\|f-g\|_{L_{2}(\mathbb{P}_{n})}^{2}$. A sub-Gaussian random variable $Z$ with mean $\rho$ and variance proxy $\tau^{2}$ can be shown to satisfy $P(Z>t+\rho)\leq\exp(-t^{2}/2\tau^{2})$ for any $t>0$ (boucheron2013, Section 2.3), and if $\rho<0$, we can take $t=-\rho$ and obtain a bound of $P(Z>0)\leq\exp(-\rho^{2}/2\tau^{2})$. Applying this result, as in the previous lemma, we conclude that conditional on $\vec{X}$ and the event $E$, $\psi=1$ holds with probability bounded by
$\exp\left(-\frac{\left(\sqrt{C^{2}-1}-2\sqrt{2}\right)^{2}\delta^{2}}{8\sigma^{2}}\right).$
The rest of the proof follows identically to Lemma 2.6. Then we can use this generalized Lemma to obtain Lemma 2.7 without any additional work, since other than using the bounds in Lemma 2.6, normality is not required. ∎
Another property we will need is the monotonicity of the local metric entropy.
Lemma 2.9.
The function $\varepsilon\mapsto M_{\mathcal{F}}^{\operatorname{loc}}(\varepsilon,c)$ is monotone non-increasing.
{toappendix}
Proof of Lemma 2.9.
We mimic the procedure in (lecamshamindramatey2022, Lemma 9). First, observe that it suffices to prove for any $h\in\mathcal{F}$ that $\varepsilon\mapsto M(\varepsilon/c,B(h,\varepsilon)\cap\mathcal{F})$ is monotone non-increasing. To see this, suppose $\varepsilon<\varepsilon^{\prime}$ and that we have proven the monotonicity for each fixed $h\in\mathcal{F}$. Suppose to the contrary that $M_{\mathcal{F}}^{\operatorname{loc}}(\varepsilon)<M_{\mathcal{F}}^{\operatorname{loc}}(\varepsilon^{\prime}).$ Then there is some $\widetilde{h}\in\mathcal{F}$ such that
$M(\varepsilon^{\prime}/c,B(\widetilde{h},\varepsilon^{\prime})\cap\mathcal{F})>M_{\mathcal{F}}^{\operatorname{loc}}(\varepsilon)=\sup_{h\in\mathcal{F}}M(\varepsilon/c,B(h,\varepsilon)\cap\mathcal{F})\geq M(\varepsilon/c,B(\widetilde{h},\varepsilon)\cap\mathcal{F}),$
violating our monotonicity property with $\widetilde{h}$.
Now fix $h\in\mathcal{F}$ and let us prove $\varepsilon\mapsto M(\varepsilon/c,B(h,\varepsilon)\cap\mathcal{F})$ is monotone non-increasing. Suppose $\varepsilon<\varepsilon^{\prime}$. Let $M=M(\varepsilon^{\prime}/c,B(h,\varepsilon^{\prime})\cap\mathcal{F})$ and let $f_{1},\dots,f_{M}$ be a maximal $\varepsilon^{\prime}/c$-packing of $B(h,\varepsilon^{\prime})\cap\mathcal{F}$. We must now construct a $\varepsilon/c$-packing of cardinality at least $M$ of the set $B(h,\varepsilon)\cap\mathcal{F}$. Our strategy will be to pull the chosen packing functions closer to $h$. Namely, we set $g_{i}=\left(1-\frac{\varepsilon}{\varepsilon^{\prime}}\right)h+\frac{\varepsilon}{\varepsilon^{\prime}}f_{i},$ noting that $g_{i}\in\mathcal{F}$ for all $i\in[M]$ by convexity of $\mathcal{F}$. Further, each $g_{i}\in B(h,\varepsilon)$ since
$\|g_{i}-h\|_{L_{2}(\mathbb{P}_{X})}$ $=\left\|\frac{\varepsilon}{\varepsilon^{\prime}}\cdot f_{i}-\frac{\varepsilon}{\varepsilon^{\prime}}\cdot h\right\|_{L_{2}(\mathbb{P}_{X})}=\frac{\varepsilon}{\varepsilon^{\prime}}\cdot\|f_{i}-h\|_{L_{2}(\mathbb{P}_{X})}\leq\frac{\varepsilon}{\varepsilon^{\prime}}\cdot\varepsilon^{\prime}=\varepsilon,$
where we used the fact that the $f_{i}\in B(h,\varepsilon^{\prime})$. Lastly, for $i\neq j$, since the $f_{i}$ are $\varepsilon^{\prime}/c$-separated,
$\|g_{i}-g_{j}\|_{L_{2}(\mathbb{P}_{X})}$ $=\left\|\frac{\varepsilon}{\varepsilon^{\prime}}(f_{i}-f_{j})\right\|_{L_{2}(\mathbb{P}_{X})}>\frac{\varepsilon}{\varepsilon^{\prime}}\cdot\frac{\varepsilon^{\prime}}{c}=\frac{\varepsilon}{c}.$
Thus $g_{1},\dots,g_{M}$ is indeed a $\varepsilon/c$-packing set of $B(h,\varepsilon)\cap\mathcal{F}$ of cardinality $M$. Hence the maximal such packing set has cardinality at least $M$. ∎
{toappendix}
The following technical lemma will be useful for the proof of Theorem 2.11.
Lemma 2.10.
For any integer $J>1$ and constant $\rho>0$, we have
$\sum_{j=2}^{J}\exp\left(-\rho\cdot\frac{1}{2^{2(j-1)}}\right)\leq\frac{\exp\left(-\rho\cdot\frac{1}{2^{2(J-1)}}\cdot\right)}{1-\exp\left(-\rho\cdot\frac{1}{2^{2(J-1)}}\cdot\right)}.$
Proof of Lemma 2.10.
Note first that $\exp\left(-\rho\cdot\frac{1}{4^{J-1}}\cdot\right)\in(0,1)$. Then
$\sum_{j=2}^{J}\exp\left(-\rho\cdot\frac{1}{2^{2(j-1)}}\right)$ $=\sum_{j=1}^{J-1}\exp\left(-\rho\cdot\frac{1}{4^{j}}\right)$
$=\sum_{j=1}^{J-1}\exp\left(-\rho\cdot\frac{1}{4^{J-1}}\cdot\frac{1}{4^{j+1-J}}\right)$
$=\sum_{j=1}^{J-1}\left[\exp\left(-\rho\cdot\frac{1}{4^{J-1}}\cdot\right)\right]^{\frac{1}{4^{j+1-J}}}$
$=\sum_{j=1}^{J-1}\left[\exp\left(-\rho\cdot\frac{1}{4^{J-1}}\cdot\right)\right]^{4^{J-1-j}}$
$=\sum_{k=0}^{J-2}\left[\exp\left(-\rho\cdot\frac{1}{4^{J-1}}\cdot\right)\right]^{4^{k}}$
$\leq\sum_{k=0}^{\infty}\left[\exp\left(-\rho\cdot\frac{1}{4^{J-1}}\cdot\right)\right]^{4^{k}}$
$\leq\sum_{k=1}^{\infty}\left[\exp\left(-\rho\cdot\frac{1}{4^{J-1}}\right)\right]^{k}$
$=\frac{\exp\left(-\rho\cdot\frac{1}{4^{J-1}}\cdot\right)}{1-\exp\left(\rho\cdot\frac{1}{4^{J-1}}\right)}.$
Now we return to our algorithm and quantify its error in the $L_{2}(\mathbb{P}_{X})$-distance. We then show in the subsequent theorem that the minimum of this error and the diameter is in fact the minimax rate. By Remark 2.8, the proof holds for sub-Gaussian noise as well.
Theorem 2.11.
Let $J^{\ast}$ be the maximal integer $J\geq 1$ with $\varepsilon_{J}=\frac{d\sqrt{L(c/2-1,\sigma,B_{\mathcal{F}})}}{2^{J-2}c}$ satisfying
$n\varepsilon_{J}^{2}>2\log M_{\mathcal{F}}^{\operatorname{loc}}\left(\frac{\varepsilon_{J}\cdot c}{\sqrt{L(c/2-1,\sigma,B_{\mathcal{F}})}},c\right)\vee\log 2,$ (2.6)
and $J^{\ast}=1$ if this condition is never satisfied. Let $f^{\ast}(\vec{X},\vec{Y})$ be the output of $J^{\ast}$ steps of Algorithm 1. Then $\mathbb{E}_{X,Y}\|\bar{f}-f^{\ast}\|_{L_{2}(\mathbb{P}_{X})}^{2}\leq\overline{C}{\varepsilon_{J^{\ast}}}^{2}$ for some constant $\overline{C}$ depending only on universal constants, $B_{\mathcal{F}}$, and $\sigma$. We defined $L(C,\sigma,B_{\mathcal{F}})$ in (2.3), where $C$ is some sufficiently large universal constant and $c:=2(C+1)$ is the constant we use in the definition of local metric entropy.
{toappendix}
Proof of Theorem 2.11.
Suppose $J>1$ steps of the algorithm are performed and consider $2\leq j\leq J$. Let $\Upsilon_{j+1}$ be the function added to $\Upsilon$ in step $j$ (e.g., in the first iteration of the loop, we append $\Upsilon_{2}$). During step $j-1$, we start with some $\Upsilon_{j-1}$ and then form a $\frac{d}{2^{j-1}(C+1)}$-packing $M_{j-1}$ of the set $B(\Upsilon_{j-1},d/2^{j-2})\cap\mathcal{F}$. We then set $\Upsilon_{j}=\mathop{\mathrm{argmin}}_{f\in M_{j-1}}\sum_{i=1}^{n}(Y_{i}-f(X_{i}))^{2}$. Applying Lemma 2.7 with $\delta=\frac{d}{2^{j-1}(C+1)}$, $\mathcal{F^{\prime}}:=B(\Upsilon_{j-1},d/2^{j-2})\cap\mathcal{F}$, and conditioning on $\|\bar{f}-\Upsilon_{j-1}\|_{L_{2}(\mathbb{P}_{X})}\leq d/2^{j-2}$ (which implies $\bar{f}\in\mathcal{F}^{\prime}$), we obtain
$\mathbb{P}_{\bar{f}}\left(\|\bar{f}-\Upsilon_{j}\|_{L_{2}(\mathbb{P}_{X})}\geq\frac{d}{2^{j-1}}\bigg{|}\|\bar{f}-\Upsilon_{j-1}\|_{L_{2}(\mathbb{P}_{X})}\leq\frac{d}{2^{j-2}},\Upsilon_{j-1}\right)\leq 3\cdot|M_{j-1}|\exp(-n\varepsilon_{j}^{2}),$
where we define
$\varepsilon_{j}$ $=\frac{d\sqrt{L(C,\sigma,B_{\mathcal{F}})}}{2^{j-1}(C+1)}.$
Further, define $a_{j}=\exp(-n\varepsilon_{j}^{2})$. Observe that $|M_{j-1}|$ is upper-bounded by $M_{\mathcal{F}}^{\operatorname{loc}}(d/2^{j-2},2(C+1))$, and by Lemma 2.9 this quantity is bounded by $M_{\mathcal{F}}^{\operatorname{loc}}(d/2^{J-2},2(C+1))$. The probability upper bound does not depend on $\Upsilon_{j-1}$, so we have proven that
$\mathbb{P}_{\bar{f}}\left(\|\bar{f}-\Upsilon_{j}\|_{L_{2}(\mathbb{P}_{X})}\geq\frac{d}{2^{j-1}}\bigg{|}\|\bar{f}-\Upsilon_{j-1}\|_{L_{2}(\mathbb{P}_{X})}\leq\frac{d}{2^{j-2}}\right)$ $\leq 3\cdot M_{\mathcal{F}}^{\operatorname{loc}}(d/2^{J-2},2(C+1))\cdot\exp(-n\varepsilon_{j}^{2}).$
If we set $A_{j}$ to be the event that $\|\bar{f}-\Upsilon_{j}\|_{L_{2}(\mathbb{P}_{X})}\geq\frac{d}{2^{j-1}}$, then we can use the property that $P(A_{J})\leq P(A_{J}|A_{J-1}^{c})+P(A_{j-1}|A_{j-2}^{c})+\dots+P(A_{3}|A_{2}^{c})+P(A_{2})$ for each $i$ to conclude that
$\mathbb{P}_{\bar{f}}\left(\|\bar{f}-\Upsilon_{J}\|_{L_{2}(\mathbb{P}_{X})}\geq\frac{d}{2^{J-1}}\right)$ $\leq 3\cdot M_{\mathcal{F}}^{\operatorname{loc}}(d/2^{J-2},2(C+1))\cdot\sum_{j=2}^{J}a_{j}$
$\leq 3\cdot\mathbbm{1}(J>1)M_{\mathcal{F}}^{\operatorname{loc}}(d/2^{J-2},2(C+1))\frac{a_{J}}{1-a_{J}}.$
The final inequality used Lemma 2.10. Note that $a_{J}<1$ since we assume $C$ is large enough that $L(C,\sigma,B_{\mathcal{F}})>0$.
Note that
$\frac{d}{2^{J-2}}=\frac{2\varepsilon_{J}(C+1)}{\sqrt{L(C,\sigma,B_{\mathcal{F}})}}.$
Now, if $n\varepsilon_{J}^{2}>2\log M_{\mathcal{F}}^{\operatorname{loc}}\left(\frac{2\cdot\varepsilon_{J}(C+1)}{\sqrt{L(C,\sigma,B_{\mathcal{F}})}},2(C+1)\right)$ and $a_{J}=\exp\left(-n\varepsilon_{J}^{2}\right)<\frac{1}{2}$, which implies $(1-a_{J})^{-1}<2$, then we have
$M_{\mathcal{F}}^{\operatorname{loc}}(d/2^{J-2},2(C+1))\cdot\frac{a_{J}}{1-a_{J}}$ $=M_{\mathcal{F}}^{\operatorname{loc}}\left(\frac{2\varepsilon_{J}(C+1)}{\sqrt{L(C,\sigma,B_{\mathcal{F}})}},2(C+1)\right)\cdot\frac{a_{J}}{1-a_{J}}$
$<\exp\left(\frac{n\varepsilon_{J}^{2}}{2}\right)\cdot\frac{a_{J}}{1-a_{J}}$
$\leq\exp\left(\frac{n\varepsilon_{J}^{2}}{2}\right)\cdot\exp\left(-n\varepsilon_{J}^{2}\right)\cdot\frac{1}{1-a_{J}}$
$\leq 2\exp\left(-\frac{n\varepsilon_{J}^{2}}{2}\right).$
Since $a_{J}<\frac{1}{2}$ iff $n\varepsilon_{J}^{2}>\log 2$, we conclude that if
$n\varepsilon_{J}^{2}>2\log M_{\mathcal{F}}^{\operatorname{loc}}\left(\frac{2\cdot\varepsilon_{J}(C+1)}{\sqrt{L(C,\sigma,B_{\mathcal{F}})}},2(C+1)\right)\vee\log 2,$
then $M_{\mathcal{F}}^{\operatorname{loc}}(d/2^{J-2},2(C+1))\cdot\frac{a_{J}}{1-a_{J}}\leq 2\exp\left(-\frac{n\varepsilon_{J}^{2}}{2}\right)$.
To summarize, if condition (2.6) holds for $\varepsilon_{J}$, then we have
$\mathbb{P}_{\bar{f}}\left(\|\bar{f}-\Upsilon_{J}\|_{L_{2}(\mathbb{P}_{X})}\geq\frac{d}{2^{J-1}}\right)$ $\leq 3\cdot\mathbbm{1}(J>1)M_{\mathcal{F}}^{\operatorname{loc}}(d/2^{J-2},2(C+1))\cdot\frac{a_{J}}{1-a_{J}}$
$\leq 6\cdot\mathbbm{1}(J>1)\exp\left(-\frac{n\varepsilon_{J}^{2}}{2}\right).$
In other words, (2.6) gives us a lower bound on $P(A_{J}^{c}).$
Now, set $\kappa=\frac{3(C+1)}{\sqrt{L(C,\sigma,B_{\mathcal{F}})}}$. Suppose $J$ satisfies condition (2.6). Then $J\leq J^{\ast}$ with $f^{\ast}$ the output of $J^{\ast}$ steps of the algorithm. By the Cauchy sequence argument in (2.2), we have $\|f^{\ast}-\Upsilon_{J}\|_{L_{2}(\mathbb{P}_{X})}\leq\frac{d}{2^{J-2}}$. Now if we assume further that $\|\Upsilon_{J}-\bar{f}\|_{L_{2}(\mathbb{P}_{X})}\leq\frac{d}{2^{J-1}}$, i.e., the event $A_{J}^{c}$, we have by the triangle inequality that $\|f^{\ast}-\bar{f}\|_{L_{2}(\mathbb{P}_{X})}\leq\frac{3d}{2^{J-1}}=\kappa\varepsilon_{J}$. If we call $B_{J}$ the event that $\|f^{\ast}-\bar{f}\|_{L_{2}(\mathbb{P}_{X})}\leq\kappa\varepsilon_{J}$, we have $P(A_{J}^{c})\leq P(B_{J})$ for $J$ satisfying (2.6). We already obtained a lower bound on $P(A_{J}^{c})$ under this condition, so in summary, for all $J$ satisfying (2.6), we have a lower bound on $P(B_{J})$, namely, that
$\|\bar{f}-f^{\ast}\|_{L_{2}(\mathbb{P}_{X})}$ $\leq\kappa\cdot\varepsilon_{J}$
with probability at least
$1-6\cdot\exp\left(-\frac{n\varepsilon_{J}^{2}}{2}\right).$
Let $J^{\ast}$ be the maximal $J$ for which conditions (2.6) all hold, with $J^{\ast}=1$ if none do. Note that $\varepsilon_{J}\downarrow 0$ as $J\to\infty$, so (2.6) will not hold for all sufficiently large $J$ (the left-side of (2.6) is increasing while the right-hand side is non-increasing). This implies $J^{\ast}<\infty$. Further, set $\underline{C}=6$, $C^{\prime}=\frac{1}{2}$. Then for any $1\leq J\leq J^{\ast}$ including $J^{\ast}=1$, we have
$\mathbb{P}_{\bar{f}}\left(\|\bar{f}-f^{\ast}\|_{L_{2}(\mathbb{P}_{X})}\geq\kappa\cdot\varepsilon_{J}\right)$ $\leq\underline{C}\cdot\mathbbm{1}(J>1)\exp(-C^{\prime}n\varepsilon_{J}^{2})$
$\leq\underline{C}\cdot\mathbbm{1}(J^{\ast}>1)\exp(-C^{\prime}n\varepsilon_{J}^{2})$ (2.7)
Note the first inequality trivially holds for $J=1$ since $\kappa\varepsilon_{1}=3d$, so that the probability will be 0 as will the right-hand side. Similarly, we extend this to hold for all $J\in\mathbb{Z}$ with $J\leq J^{\ast}$, as if $J\leq 0$, then $\kappa\cdot\varepsilon_{J}=\frac{3d}{2^{J-1}}\geq 6d$. But since $d$ is the diameter of $\mathcal{F}$, the above probability is 00 which is less than than the right hand side of (2.7).
Now observe that $\bigcup_{J\leq J^{\ast}}[\varepsilon_{J},\varepsilon_{J-1})=[\varepsilon_{J^{\ast}},\infty)$. Pick any $x\geq\varepsilon_{J^{\ast}}$. Then $x\in[\varepsilon_{J},\varepsilon_{J-1})$ for some $J\leq J^{\ast}$. Thus
$2\kappa x\geq 2\kappa\cdot\varepsilon_{J}=\kappa\cdot\varepsilon_{J-1}.$
Note further that $x\mapsto\exp(-C^{\prime}nx^{2})$ is non-increasing for $x\geq 0$. Consequently,
$\mathbb{P}_{\bar{f}}\left(\|\bar{f}-f^{\ast}\|_{L_{2}(\mathbb{P}_{X})}\geq 2\kappa x\right)$ $\leq\mathbb{P}_{\bar{f}}\left(\|\bar{f}-f^{\ast}\|_{L_{2}(\mathbb{P}_{X})}\geq\kappa\cdot\varepsilon_{J-1}\right)$
$\leq\underline{C}\cdot\mathbbm{1}(J^{\ast}>1)\exp(-C^{\prime}n\varepsilon_{J-1}^{2})$
$\leq\underline{C}\cdot\mathbbm{1}(J^{\ast}>1)\exp(-C^{\prime}nx^{2}).$
This proves that $\mathbb{P}_{\bar{f}}\left(\|\bar{f}-f^{\ast}\|_{L_{2}(\mathbb{P}_{X})}\geq 2\kappa x\right)\leq\underline{C}\cdot\mathbbm{1}(J^{\ast}>1)\exp(-C^{\prime}nx^{2})$ for all $x\geq\varepsilon_{J^{\ast}}$.
Then we compute $\mathbb{E}_{X,Y}\|\bar{f}-f^{\ast}\|_{L_{2}(\mathbb{P}_{X})}^{2}$ by integration, splitting the integral into two parts:
$\mathbb{E}_{X,Y}\|\bar{f}-f^{\ast}\|_{L_{2}(\mathbb{P}_{X})}^{2}$ $=\int_{0}^{\infty}2x\mathbb{P}_{\bar{f}}\left(\|\bar{f}-f^{\ast}\|_{L_{2}(\mathbb{P}_{X})}\geq x\right)\mathrm{d}x$
$\leq\int_{0}^{2\kappa\varepsilon_{J^{\ast}}}2x\mathrm{d}x+\int_{2\kappa\varepsilon_{J^{\ast}}}^{\infty}2x\mathbb{P}_{\bar{f}}\left(\|\bar{f}-f^{\ast}\|_{L_{2}(\mathbb{P}_{X})}\geq x\right)\mathrm{d}x$
$\leq C^{\prime\prime\prime}{\varepsilon_{J^{\ast}}}^{2}+\int_{2\kappa\varepsilon_{J^{\ast}}}^{\infty}2x\cdot\underline{C}\cdot\mathbbm{1}(J^{\ast}>1)\exp(-C^{\prime\prime}nx^{2})\mathrm{d}x$
$=C^{\prime\prime\prime}{\varepsilon_{J^{\ast}}}^{2}+C^{\prime\prime\prime\prime}\cdot\mathbbm{1}(J^{\ast}>1)n^{-1}\exp(-C^{\prime\prime\prime\prime\prime}n{\varepsilon_{J^{\ast}}}^{2}).$
Recall from our assumption (2.6) that either $J^{\ast}=1$ or $n{\varepsilon_{J^{\ast}}}^{2}>\log 2$. In either case, $\mathbb{E}_{X,Y}\|\bar{f}-f^{\ast}\|_{L_{2}(\mathbb{P}_{X})}^{2}\leq\overline{C}{\varepsilon_{J^{\ast}}}^{2}$ for some constant $\overline{C}$ depending only on $\sigma$, $B_{\mathcal{F}}$, and our constant $C>3$.
Theorem 2.12.
Define $\varepsilon^{\ast}:=\sup\{\varepsilon:n\varepsilon^{2}\leq\log M_{\mathcal{F}}^{\operatorname{loc}}(\varepsilon,c)\}$ where $c$ is a sufficiently large absolute constant. Then the minimax rate is ${\varepsilon^{\ast}}^{2}\wedge d^{2}$ up to absolute constants.
{toappendix}
Proof of Theorem 2.12.
Suppose $n{\varepsilon^{\ast}}^{2}>8\log 2$. Set $\alpha=\min(\sigma/2,1)\in(0,1]$ and $\delta^{\ast}=\alpha\varepsilon^{\ast}$. Note that $\alpha^{2}\leq\frac{\sigma^{2}}{4}$. Then using the monotonicity of $M_{\mathcal{F}}^{\operatorname{loc}}(\cdot,c)$, we obtain
$\log M_{\mathcal{F}}^{\operatorname{loc}}(\delta^{\ast},c)$ $\geq\log M_{\mathcal{F}}^{\operatorname{loc}}(\varepsilon^{\ast},c)\geq n{\varepsilon^{\ast}}^{2}=\frac{n{\varepsilon^{\ast}}^{2}}{2}+\frac{n{\varepsilon^{\ast}}^{2}}{2}>\frac{n{\delta^{\ast}}^{2}}{2\alpha^{2}}+4\log 2$
$\geq 4\cdot\frac{n{\delta^{\ast}}^{2}}{2\sigma^{2}}+4\log 2\geq 4\left(\frac{n{\delta^{\ast}}^{2}}{2\sigma^{2}}\vee\log 2\right).$
Thus, the lower bound condition from Lemma 2.3 is satisfied for $\delta^{\ast}$, so indeed ${\varepsilon^{\ast}}^{2}$ is a lower bound for the minimax rate up to absolute constants by the lemma when $n{\varepsilon^{\ast}}^{2}>8\log 2$.
We now show the minimax rate has an upper bound of rate ${\varepsilon^{\ast}}^{2}$ when $n{\varepsilon^{\ast}}^{2}>8\log 2$. Following lecamshamindramatey2022, we set $\eta=\frac{c\sqrt{2}}{L(c/2-1,\sigma,B_{\mathcal{F}})}\wedge 1\in(0,1)$, pick $D$ sufficiently large such that $D\eta>1$, and set $\delta=D\sqrt{2}\varepsilon^{\ast}$. Again using monotonicity of $M_{\mathcal{F}}^{\operatorname{loc}}(\cdot,c)$ and the definition of $\varepsilon^{\ast}$, we have
$n\delta^{2}$ $=2D^{2}n{\varepsilon^{\ast}}^{2}=\frac{2}{\eta^{2}}\cdot n(D\eta\cdot\varepsilon^{\ast})^{2}>\frac{2}{\eta^{2}}\cdot\log M_{\mathcal{F}}^{\operatorname{loc}}(D\eta\varepsilon^{\ast},c)\geq 2\log M_{\mathcal{F}}^{\operatorname{loc}}(D\eta\varepsilon^{\ast},c)$
$\geq 2\log M_{\mathcal{F}}^{\operatorname{loc}}\left(\frac{c\sqrt{2}\cdot D\varepsilon^{\ast}}{L(c/2-1,\sigma,B_{\mathcal{F}})},c\right)=2\log M_{\mathcal{F}}^{\operatorname{loc}}\left(\frac{c\cdot\delta}{L(c/2-1,\sigma,B_{\mathcal{F}})},c\right)$
Further, since $n{\varepsilon^{\ast}}^{2}>8\log 2$, we have $n\delta^{2}=2nD^{2}{\varepsilon^{\ast}}^{2}>16D^{2}\log 2>\log 2$, so indeed
$n\delta^{2}>2\log M_{\mathcal{F}}^{\operatorname{loc}}\left(\frac{\delta\cdot c}{\sqrt{L(c/2-1,\sigma,B_{\mathcal{F}})}},c\right)\vee\log 2.$
In other words, $\delta$ satisfies the condition in (2.6).
Now consider the map
$\psi\colon 0<x\mapsto nx^{2}-2\log M_{\mathcal{F}}^{\operatorname{loc}}\left(x\cdot\frac{c}{\sqrt{L(c/2-1,\sigma,B_{\mathcal{F}})}},c\right)\vee\log 2$
which is non-decreasing. We have shown that $\psi(\delta)>0$. Recall our definition of $\varepsilon_{J^{\ast}}$ from Theorem 2.11, and let us show that $\delta\geq\varepsilon_{J^{\ast}}/2$. Suppose no $J\geq 1$ exists such that $\varepsilon_{J}$ satisfies (2.6), in which case $J^{\ast}=1$. Then $\psi(\varepsilon_{J^{\ast}}/2)\leq\psi(\varepsilon_{J^{\ast}})<0<\psi(\delta)$ so $\delta\geq\varepsilon_{J^{\ast}}/2$. On the other hand, suppose some $J$ does exist such that $\varepsilon_{J}$ satisfies (2.6). Because $J^{\ast}<\infty$ is chosen maximally in our algorithm, $\varepsilon_{J^{\ast}+1}=\varepsilon_{J^{\ast}}/2$ does not satisfy (2.6), so we have $\psi(\varepsilon_{J^{\ast}}/2)=\psi(\varepsilon_{J^{\ast}+1})<0<\psi(\delta)$, so $\delta\geq\varepsilon_{J^{\ast}}/2$.
Since the minimax rate has an upper bound of $\varepsilon_{J^{\ast}}^{2}$ up to constants by Theorem 2.11, it follows that the rate is also upper bounded by $4\delta^{2}=8D^{2}{\varepsilon^{\ast}}^{2}$. Thus, the minimax has a lower and upper bound of rate ${\varepsilon^{\ast}}^{2}$ up to absolute constants. But if ${\varepsilon^{\ast}}^{2}$ is a lower bound, then so is ${\varepsilon^{\ast}}^{2}\wedge d$; further, the squared diameter is always an upper bound for the minimax rate, so ${\varepsilon^{\ast}}^{2}\wedge d^{2}$ is also an upper bound up to absolute constants. Thus, if $n{\varepsilon^{\ast}}^{2}>8\log 2$, the minimax rate is ${\varepsilon^{\ast}}^{2}\wedge d^{2}$ up to absolute constants.
Next, we consider the case where $n{\varepsilon^{\ast}}^{2}\leq 8\log 2$. Observe then that
$\log M_{\mathcal{F}}^{\operatorname{loc}}(2\varepsilon^{\ast},c)\leq 4n{\varepsilon^{\ast}}^{2}\leq 32\log 2.$
Now take two endpoints of a diameter of $\mathcal{F}$ and form a ball $B$ of radius $2\varepsilon^{\ast}$ centered at the midpoint of these two endpoints. Consider the diameter of $B$ that coincides with the chosen diameter of $\mathcal{F}$ at infinitely many points. Since an interval of length $4\varepsilon^{\ast}$ if partitioned into subintervals of length less than $(2\varepsilon^{\ast}/c)$ is composed of at least $2c$ subintervals, for sufficiently large $c$, at least $\exp(32\log 2)$ many points at least $(2\varepsilon^{\ast}/c)$ distance apart can be chosen along this particular diameter of $B$. All of these points cannot also belong to $\mathcal{F}$, otherwise we will have formed a $(2\varepsilon^{\ast}/c)$-packing of $B\cap\mathcal{F}$ with $\exp(32\log 2)$ points, contradicting $\log M_{\mathcal{F}}^{\operatorname{loc}}(2\varepsilon^{\ast},c)\leq 32\log 2$. This means the endpoints of the diameter of $\mathcal{F}$ are strictly contained in the diameter of $B$ we picked, so that $d\leq 4\varepsilon^{\ast}\leq 8\sqrt{2n^{-1}\log 2}$.
Now, consider an $\varepsilon$ which is proportional to $d$. In this case, $n\varepsilon^{2}$ is upper bounded by some constant since $d\leq 8\sqrt{2n^{-1}\log 2}$, while $\log M_{\mathcal{F}}^{\operatorname{loc}}(2\varepsilon,c)$ can be made arbitrarily large by increasing $c$, to ensure the condition from Lemma 2.3 for the lower bound is attained. So $\varepsilon^{2}$ is a lower bound for the minimax-rate up to constants, and thus by proportionality, $d^{2}$ is a lower bound up to absolute constants. Thus ${\varepsilon^{\ast}}^{2}\wedge d^{2}$ is a lower bound for the minimax rate.
But $d^{2}$ is also always an upper bound for the minimax rate, so the minimax rate is of order $d^{2}$ up to constants. But in this second case where $n{\varepsilon^{\ast}}^{2}\leq 8\log 2$, we have $d\leq 4\varepsilon^{\ast}$, showing that ${\varepsilon^{\ast}}^{2}$ and thus ${\varepsilon^{\ast}}^{2}\wedge d^{2}$ is also an upper bound up to absolute constants. We conclude therefore the minimax rate is ${\varepsilon^{\ast}}^{2}\wedge d^{2}$ up to constants. ∎
Remark 2.13 (Without Total Boundedness).
Consider the case where the local packing number is infinite for some $\varepsilon$, i.e., there are arbitrarily large $\varepsilon/c$-packings of sets of the form $\mathcal{F}\cap B(f,\varepsilon)$ for $f\in\mathcal{F}$. Note that Lemma 2.1 and Lemma 2.2 still hold since they only require the existence of $\varepsilon$-separated sets. For the lower bound result in Lemma 2.3, we adapt the proof by picking an $f$ such that the maximal $(\varepsilon/c)$-packing set of $\mathcal{F}\cap B(f,\varepsilon)$ is infinite. Then we use a finite subset of that infinite packing set and apply the same logic as in the rest of the proof.
For the upper bound, the algorithm may require solving an optimization over infinitely many functions, but our results still hold. Lemma 2.5 and Lemma 2.6 do not rely on any packing constructions but simply a concentration result. Lemma 2.7 is a statement about any packing set, not necessarily a maximal or infinite one. Monotonicity of the local metric entropy is unaffected. The arguments within the proof of Theorem 2.11 and Theorem 2.12 also hold either trivially (e.g., that the probability is bounded by infinity) or a key hypothesis such as (2.6) implies the local metric entropy is finite.
The similar proofs in the following unbounded and adaptive settings will also be unaffected.
3 Extension to the Unbounded Case
Let us now assume $\mathcal{F}$ is not bounded in sup-norm (i.e., for all $r>0$, there is an $f\in\mathcal{F}$ and $x\in\mathcal{X}$ such that $|f(x)|>r$). But recall that we assume the diameter $d:=\mathrm{diam}(\mathcal{F}):=\sup\{\|f-g\|_{L_{2}(\mathbb{P}_{X})}:f,g\in\mathcal{F}\}$ is bounded.
However, we require an additional assumption similar to mendelson_2017_local_versus_global, namely, that there is some $\alpha>0$ such that for each $p\geq 2$ and $f,g\in\mathcal{F}$ we have
$\|f-g\|_{L_{p}(\mathbb{P}_{X})}\leq\alpha\sqrt{p}\|f-g\|_{L_{2}(\mathbb{P}_{X})}.$ (3.1)
Here we define for any $p\geq 2$ the $L_{p}(\mathbb{P}_{X})$-distance between $f,g\in\mathcal{F}$ by
$\|f-g\|_{L_{p}(\mathbb{P}_{X})}=\left(\int|f(x)-g(x)|^{p}\mathrm{d}\mathbb{P}_{X}(x)\right)^{1/p}.$
We omit the assumption about boundedness of the induced centered canonical Gaussian process $\{G_{f}:f\in\mathcal{F}\}$ used in mendelson_2017_local_versus_global. Rather, we assume boundedness of the diameter. Moreover, our approach is able to match our lower and upper bounds.
Our lower bounds from Section 2 all carry through with no change. For the upper bound, we will need slightly different concentration inequalities, but several of the proofs will carry over with minimal changes.
Define the subexponential norm $\|\cdot\|_{\psi_{1}}$ of a random variable $Z$ by $\|Z\|_{\psi_{1}}=\sup_{p\geq 1}p^{-1}\cdot\left(\mathbb{E}|Z|^{p}\right)^{1/p}.$ Then using Minkowski’s inequality and (3.1), we have
$\left|(f(X)-g(X))^{2}-\|f-g\|_{L_{2}(\mathbb{P}_{X})}^{2}\right|_{\psi_{1}}$ $=\sup_{p\geq 1}p^{-1}\sqrt[p]{\mathbb{E}|(f(X)-g(X))^{2}-\|f-g\|_{L_{2}(\mathbb{P}_{X})}^{2}|^{p}}$
$\leq\sup_{p\geq 1}p^{-1}\left[\sqrt[p]{\mathbb{E}|f(X)-g(X)|^{2p}}+\sqrt[p]{\|f-g\|_{L_{2}(\mathbb{P}_{X})}^{2p}}\right]$
$=\sup_{p\geq 1}p^{-1}\left[\left(\sqrt[2p]{\mathbb{E}|f(X)-g(X)|^{2p}}\right)^{2}+\|f-g\|_{L_{2}(\mathbb{P}_{X})}^{2}\right]$
$=\sup_{p\geq 1}p^{-1}\left[\|f-g\|_{L_{2p}(\mathbb{P}_{X})}^{2}+\|f-g\|_{L_{2}(\mathbb{P}_{X})}^{2}\right]$
$\leq\sup_{p\geq 1}p^{-1}\left(2\alpha^{2}p+1\right)\|f-g\|_{L_{2}(\mathbb{P}_{X})}^{2}$
$\leq(2\alpha^{2}+1)\|f-g\|_{L_{2}(\mathbb{P}_{X})}^{2}$
$\leq(2\alpha^{2}+1)d\|f-g\|_{L_{2}(\mathbb{P}_{X})}$
$\leq(2\alpha^{2}+1)d^{2}.$ (3.2)
We now use this result along with concentration inequalities for subexponential random variables from vershynin_2012 to obtain the following analogue of Lemma 2.5. As before, the result holds when the noise is sub-Gaussian.
Lemma 3.1.
Suppose $\mathcal{F}$ satisfies (3.1), and $f,g\in\mathcal{F}$ satisfy $n\|f-g\|_{L_{2}(\mathbb{P}_{X})}^{2}\geq C^{2}\delta^{2}$ for some sufficiently large $C>0$. Suppose further that $n\|f-\bar{f}\|_{L_{2}(\mathbb{P}_{X})}^{2}<\delta^{2}$. Then $n\|f-\bar{f}\|_{L_{2}(\mathbb{P}_{n})}^{2}\leq 2\delta^{2}$ and $n\|f-g\|_{L_{2}(\mathbb{P}_{n})}^{2}\geq(C^{2}-1)\delta^{2}$ both simultaneously hold with probability at least
$1-2\exp\left(-\frac{b\delta^{2}}{C^{2}(2\alpha^{2}+1)^{2}d^{2}}\right),$
where $b>0$ is some universal constant.
{toappendix}
Proof of Lemma 3.1.
Consider the random variables $Z_{i}=(f(X_{i})-g(X_{i}))^{2}-\|f-g\|_{L_{2}(\mathbb{P}_{X})}^{2}$, which are centered and sub-exponential (vershynin_2012, Definition 5.13 or property (5.15)) with $\|Z_{i}\|_{\psi_{1}}\leq(2\alpha^{2}+1)d^{2}$ for each $i$ by (3.2). Set $K:=\max_{1\leq i\leq n}\|Z_{i}\|_{\psi_{1}}$. By Proposition 5.16 of vershynin_2012, we have for some absolute constant $b>0$ that
$\mathbb{P}_{\bar{f}}\left(\sum_{i=1}^{n}Z_{i}<-t\right)$ $\leq\exp\left(-b\cdot\min\left(\frac{t^{2}}{nK^{2}},\frac{t}{K}\right)\right).$
Then setting $t=C^{-2}n\|f-g\|_{L_{2}(\mathbb{P}_{X})}^{2}$,
$\mathbb{P}_{\bar{f}}\left(n\|f-g\|_{L_{2}(\mathbb{P}_{n})}^{2}<(C^{2}-1)\delta^{2}\right)$ $=\mathbb{P}_{\bar{f}}\left(n\|f-g\|_{L_{2}(\mathbb{P}_{n})}^{2}<(1-C^{-2})\cdot C^{2}\delta^{2}\right)$
$\leq\mathbb{P}_{\bar{f}}\left(n\|f-g\|_{L_{2}(\mathbb{P}_{n})}^{2}<(1-C^{-2})\cdot n\|f-g\|_{L_{2}(\mathbb{P}_{X})}^{2}\right)$
$=\mathbb{P}_{\bar{f}}\left(n\|f-g\|_{L_{2}(\mathbb{P}_{n})}^{2}-n\|f-g\|_{L_{2}(\mathbb{P}_{X})}^{2}<\right.$
$\quad\quad\quad\quad\quad\left.-C^{-2}\cdot n\|f-g\|_{L_{2}(\mathbb{P}_{X})}^{2}\right)$
$=\mathbb{P}_{\bar{f}}\left(\sum_{i=1}^{n}Z_{i}<-t\right)$
$\leq\exp\left(-b\cdot\min\left(\frac{t^{2}}{nK^{2}},\frac{t}{K}\right)\right).$
Observe that since $K\leq(2\alpha^{2}+1)d\|f-g\|_{L_{2}(\mathbb{P}_{X})}$ (see (3.2)),
$\frac{t^{2}}{nK^{2}}$ $=\frac{C^{-4}n^{2}\|f-g\|_{L_{2}(\mathbb{P}_{X})}^{4}}{nK^{2}}\geq\frac{C^{-4}n\|f-g\|_{L_{2}(\mathbb{P}_{X})}^{4}}{(2\alpha^{2}+1)^{2}d^{2}\|f-g\|_{L_{2}(\mathbb{P}_{X})}^{2}}=\frac{C^{-4}n\|f-g\|_{L_{2}(\mathbb{P}_{X})}^{2}}{(2\alpha^{2}+1)^{2}d^{2}}$
$\geq\frac{C^{2}\delta^{2}}{C^{4}(2\alpha^{2}+1)^{2}d^{2}}=\frac{\delta^{2}}{C^{2}(2\alpha^{2}+1)^{2}d^{2}}.$
Similarly, since $K\leq(2\alpha^{2}+1)d^{2}$,
$\frac{t}{K}$ $=\frac{C^{-2}n\|f-g\|_{L_{2}(\mathbb{P}_{X})}^{2}}{K}\geq\frac{C^{-2}\cdot C^{2}\delta^{2}}{(2\alpha^{2}+1)d^{2}}=\frac{\delta^{2}}{(2\alpha^{2}+1)d^{2}}\geq\frac{\delta^{2}}{C^{2}(2\alpha^{2}+1)^{2}d^{2}},$
so $\min\left(\frac{t^{2}}{nK^{2}},\frac{t}{K}\right)\geq\frac{\delta^{2}}{C^{2}(2\alpha^{2}+1)^{2}d^{2}}$. Therefore,
$\mathbb{P}_{\bar{f}}\left(n\|f-g\|_{L_{2}(\mathbb{P}_{n})}^{2}<(C^{2}-1)\delta^{2}\right)$ $\leq\exp\left(-\frac{b\delta^{2}}{C^{2}(2\alpha^{2}+1)^{2}d^{2}}\right).$
Next, we consider the random variables $Z_{i}=(f(X_{i})-\bar{f}(X_{i}))^{2}-\|f-\bar{f}\|_{L_{2}(\mathbb{P}_{X})}^{2}$, which are again centered and sub-exponential with $\|Z_{i}\|_{\psi_{1}}\leq(2\alpha^{2}+1)d^{2}$ for each $i$. Let $K$ be defined as before. By Proposition 5.16 of vershynin_2012, we have for the same absolute constant $b>0$ that
$\mathbb{P}_{\bar{f}}\left(\sum_{i=1}^{n}Z_{i}>t\right)$ $\leq\exp\left(-b\cdot\min\left(\frac{t^{2}}{nK^{2}},\frac{t}{K}\right)\right).$
We set $t=\delta^{2}$ and obtain
$\mathbb{P}_{\bar{f}}\left(n\|f-\bar{f}\|_{L_{2}(\mathbb{P}_{n})}^{2}>2\delta^{2}\right)$ $\leq\mathbb{P}_{\bar{f}}\left(n\|f-\bar{f}\|_{L_{2}(\mathbb{P}_{n})}^{2}>n\|f-\bar{f}\|_{L_{2}(\mathbb{P}_{X})}^{2}+\delta^{2}\right)$
$=\mathbb{P}_{\bar{f}}\left(\sum_{i=1}^{n}Z_{i}>t\right)$
$\leq\exp\left(-b\cdot\min\left(\frac{t^{2}}{nK^{2}},\frac{t}{K}\right)\right).$
Since $\|f-\bar{f}\|_{L_{2}(\mathbb{P}_{X})}\leq\delta/\sqrt{n}$, we have
$K\leq(2\alpha^{2}+1)d\|f-\bar{f}\|_{L_{2}(\mathbb{P}_{X})}\leq\frac{(2\alpha^{2}+1)d\delta}{\sqrt{n}}.$
Thus,
$\frac{t^{2}}{nK^{2}}$ $=\frac{\delta^{4}}{nK^{2}}\geq\frac{n\delta^{4}}{n(2\alpha^{2}+1)^{2}d^{2}\delta^{2}}=\frac{\delta^{2}}{(2\alpha^{2}+1)^{2}d^{2}}.$
Further, since $K\leq(2\alpha^{2}+1)d^{2}$,
$\frac{t}{K}$ $=\frac{\delta^{2}}{K}\geq\frac{\delta^{2}}{(2\alpha^{2}+1)d^{2}}\geq\frac{\delta^{2}}{(2\alpha^{2}+1)^{2}d^{2}}.$
Hence $\min\left(\frac{t^{2}}{nK^{2}},\frac{t}{K}\right)\geq\frac{\delta^{2}}{(2\alpha^{2}+1)^{2}d^{2}}\geq\frac{\delta^{2}}{C^{2}(2\alpha^{2}+1)^{2}d^{2}}$. Therefore,
$\mathbb{P}_{\bar{f}}\left(n\|f-\bar{f}\|_{L_{2}(\mathbb{P}_{n})}^{2}>2\delta^{2}\right)$ $\leq\exp\left(-\frac{b\delta^{2}}{C^{2}(2\alpha^{2}+1)^{2}d^{2}}\right).$
Combining these results with a union bound, we conclude that with probability at least
$1-2\exp\left(-\frac{b\delta^{2}}{C^{2}(2\alpha^{2}+1)^{2}d^{2}}\right),$
we have $n\|f-\bar{f}\|_{L_{2}(\mathbb{P}_{n})}^{2}\leq 2\delta^{2}$ and $n\|f-g\|_{L_{2}(\mathbb{P}_{n})}^{2}\geq(C^{2}-1)\delta^{2}$. ∎
We then obtain analogues of Lemmas 2.6 and 2.7.
Lemma 3.2.
Suppose $\mathcal{F}$ satisfies (3.1), and suppose we are testing $H_{0}:\bar{f}=f$ vs $H_{1}:\bar{f}=g$ for $n\|f-g\|_{L_{2}(\mathbb{P}_{X})}^{2}\geq C^{2}\delta^{2}$ for some $C>3$. Then for some universal constant $b>0$, the test $\psi(\vec{Y})=\mathbbm{1}\left(\|\vec{Y}-f(\vec{X})\|_{2}^{2}\geq\|\vec{Y}-g(\vec{X})\|_{2}^{2}\right)$ satisfies
$\sup_{\bar{f}:n\|f-\bar{f}\|_{L_{2}(\mathbb{P}_{X})}^{2}<\delta^{2}}\mathbb{P}_{\bar{f}}(\psi=1)$ $\vee\sup_{\bar{f}:n\|g-\bar{f}\|_{L_{2}(\mathbb{P}_{X})}^{2}<\delta^{2}}\mathbb{P}_{\bar{f}}(\psi=0)\leq 3\exp\left(-\widetilde{L}(C,\sigma)\cdot\delta^{2}\right),$
where
$\widetilde{L}(C,\sigma)=\min\left\{\frac{\left(\sqrt{C^{2}-1}-2\sqrt{2}\right)^{2}}{8\sigma^{2}},\frac{b}{C^{2}(2\alpha^{2}+1)^{2}d^{2}}\right\}.$ (3.3)
{toappendix}
Proof of Lemma 3.2.
Many of the steps are identical to those from the proof of Lemma 2.6, so we omit some details. The main difference is in the specific probability bounds we obtain. Let $\eta(\vec{X})=\bar{f}(\vec{X})-f(\vec{X})$. Then as detailed in Lemma 2.6, we derive
$\|\vec{Y}-f(\vec{X})\|_{2}^{2}-\|\vec{Y}-g(\vec{X})\|_{2}^{2}$ $=-\|f(\vec{X})-g(\vec{X})\|_{2}^{2}+2\eta(\vec{X})^{T}(g(\vec{X})-f(\vec{X}))$
$\quad\quad\quad+2\xi^{T}(g(\vec{X})-f(\vec{X})).$
Observe that conditional on $\vec{X}$, by applying the Cauchy-Schwarz inequality, this quantity is normal with mean at most
$-\|f(\vec{X})-g(\vec{X})\|_{2}^{2}+2\|\eta(\vec{X})\|_{2}\|f(\vec{X})-g(\vec{X})\|_{2}$
and variance $4n\sigma^{2}\|f-g\|_{L_{2}(\mathbb{P}_{n})}^{2}$.
We now derive estimates on $\|\eta(\vec{X})\|_{2}$ and $\|f-g\|_{L_{2}(\mathbb{P}_{n})}^{2}$ that hold with high probability. Suppose $n\|f-\bar{f}\|_{L_{2}(\mathbb{P}_{X})}^{2}<\delta^{2}$. From Lemma 3.1, this condition along with our assumption $n\|f-g\|_{L_{2}(\mathbb{P}_{X})}^{2}\geq C^{2}\delta^{2}$ implies that with probability at least $1-2\exp\left(-\frac{b\delta^{2}}{C^{2}(2\alpha^{2}+1)^{2}d^{2}}\right)$, we have
$n\|f-g\|_{L_{2}(\mathbb{P}_{n})}^{2}\geq(C^{2}-1)\delta^{2},\quad n\|f-\bar{f}\|_{L_{2}(\mathbb{P}_{n})}^{2}\leq 2\delta^{2}.$ (3.4)
Let $E$ denote the event that both inequalities of (3.4) hold, so
$\mathbb{P}_{\bar{f}}(E^{c})\leq 2\exp\left(-\frac{b\delta^{2}}{C^{2}(2\alpha^{2}+1)^{2}d^{2}}\right).$
Note that the inequalities in (3.4) are the same two inequalities as in the proof of Lemma 2.6.
Hence following the details in Lemma 2.6, we obtain that
$\mathbb{P}_{\bar{f}}(\psi=1|E)\leq\exp\left(-\frac{\left(\sqrt{C^{2}-1}-2\sqrt{2}\right)^{2}\delta^{2}}{8\sigma^{2}}\right).$
Since $\mathbb{P}_{\bar{f}}(\psi=1|\vec{X},E)$ has an upper bound not depending on $X$, it follows (by taking an expectation) that
$\mathbb{P}_{\bar{f}}(\psi=1|E)\leq\exp\left(-\frac{\left(\sqrt{C^{2}-1}-2\sqrt{2}\right)^{2}\delta^{2}}{8\sigma^{2}}\right).$
Thus,
$\mathbb{P}_{\bar{f}}(\psi=1)$ $=\mathbb{P}_{\bar{f}}(\psi=1|E)\cdot\mathbb{P}_{\bar{f}}(E)+\mathbb{P}_{\bar{f}}(\psi=1|E^{c})\cdot\mathbb{P}_{\bar{f}}(E^{c})$
$\leq\mathbb{P}_{\bar{f}}(\psi=1|E)+\mathbb{P}_{\bar{f}}(E^{c})$
$\leq\exp\left(-\frac{\left(\sqrt{C^{2}-1}-2\sqrt{2}\right)^{2}\delta^{2}}{8\sigma^{2}}\right)+2\exp\left(-\frac{b\delta^{2}}{C^{2}(2\alpha^{2}+1)^{2}d^{2}}\right).$
If we instead assumed $n\|g-\bar{f}\|_{L_{2}(\mathbb{P}_{X})}^{2}<\delta^{2}$, we would arrive on an identical bound for $\mathbb{P}_{\bar{f}}(\psi=0)$. Thus, we conclude that
$\sup_{\bar{f}:n\|f-\bar{f}\|_{L_{2}(\mathbb{P}_{X})}^{2}<\delta^{2}}\mathbb{P}_{\bar{f}}(\psi=1)$ $\vee\sup_{\bar{f}:n\|g-\bar{f}\|_{L_{2}(\mathbb{P}_{X})}^{2}<\delta^{2}}\mathbb{P}_{\bar{f}}(\psi=0)\leq$
$\quad\exp\left(-\frac{\left(\sqrt{C^{2}-1}-2\sqrt{2}\right)^{2}\delta^{2}}{8\sigma^{2}}\right)+2\exp\left(-\frac{b\delta^{2}}{C^{2}(2\alpha^{2}+1)^{2}d^{2}}\right).$
We define
$\widetilde{L}(C,\sigma)=\min\left\{\frac{\left(\sqrt{C^{2}-1}-2\sqrt{2}\right)^{2}}{8\sigma^{2}},\frac{b}{C^{2}(2\alpha^{2}+1)^{2}d^{2}}\right\},$
to conclude that
$\sup_{\bar{f}:n\|f-\bar{f}\|_{L_{2}(\mathbb{P}_{X})}^{2}<\delta^{2}}\mathbb{P}_{\bar{f}}(\psi=1)$ $\vee\sup_{\bar{f}:n\|g-\bar{f}\|_{L_{2}(\mathbb{P}_{X})}^{2}<\delta^{2}}\mathbb{P}_{\bar{f}}(\psi=0)\leq 3\exp\left(-\widetilde{L}(C,\sigma)\cdot\delta^{2}\right).$
Lemma 3.3.
Suppose $\mathcal{F}$ satisfies (3.1). Let $f_{1},\dots,f_{M}$ be a $\delta$-packing of $\mathcal{F}^{\prime}\subseteq\mathcal{F}$ and suppose $\bar{f}\in\mathcal{F}^{\prime}$. Let $i^{\ast}=\mathop{\mathrm{argmin}}_{i\in[M]}\sum_{j=1}^{n}(Y_{j}-f_{i}(X_{j}))^{2}.$ Then
$\mathbb{P}_{\bar{f}}\left(\|f_{i^{\ast}}-\bar{f}\|_{L_{2}(\mathbb{P}_{X})}>\delta(C+1)\right)$ $\leq 3M\exp\left(-\widetilde{L}(C,\sigma)\cdot\delta^{2}\right),$
where $\widetilde{L}(C,\sigma)$ is defined as in (3.3).
{toappendix}
Proof of Lemma 3.3.
The proof is completely identical to that of Lemma 2.7, except in the step (2.5), we instead use Lemma 3.2. This replaces the $L(C,\sigma,B_{\mathcal{F}})$ from the bounded case with $\widetilde{L}(C,\sigma)$ where $\widetilde{L}(C,\sigma)$ is defined in (3.3).
Remark 3.4.
Lemmas 3.2 and 3.3 can be generalized to our sub-Gaussian noise setting where each $\xi_{i}\sim\mathrm{SG}(0,\sigma^{2})$. We supply the details in the appendix.
{toappendix}
Proof of Remark 3.4.
As in Lemma 3.2, we show that
$\|\vec{Y}-f(\vec{X})\|_{2}^{2}-\|\vec{Y}-g(\vec{X})\|_{2}^{2}$ $=-\|f(\vec{X})-g(\vec{X})\|_{2}^{2}+2\eta(\vec{X})^{T}(g(\vec{X})-f(\vec{X}))$
$\quad\quad\quad+2\xi^{T}(g(\vec{X})-f(\vec{X})).$
Conditional on $\vec{X}$, this quantity is sub-Gaussian with mean bounded by
$-\|f(\vec{X})-g(\vec{X})\|_{2}^{2}+2\|\eta(\vec{X})\|_{2}\|f(\vec{X})-g(\vec{X})\|_{2},$
and has variance proxy $4n\sigma^{2}\|f-g\|_{L_{2}(\mathbb{P}_{n})}^{2}$. The proof of Lemma 3.2 then defines the event $E$ as when the inequalities in (3.4) hold, and $E$ satisfies
$\mathbb{P}_{\bar{f}}(E^{c})\leq 2\exp\left(-\frac{b\delta^{2}}{C^{2}(2\alpha^{2}+1)^{2}d^{2}}\right).$
Noting that Lemma 3.1 still holds with sub-Gaussian noise, we obtain as in Lemma 3.2 that conditional on $\vec{X}$, the event $E$ implies $\|\vec{Y}-f(\vec{X})\|_{2}^{2}-\|\vec{Y}-g(\vec{X})\|_{2}^{2}$ is sub-Gaussian with mean bounded above by the negative quantity
$\left(-1+\frac{2\sqrt{2}}{\sqrt{C^{2}-1}}\right)\cdot n\|f-g\|_{L_{2}(\mathbb{P}_{n})}^{2}$
and variance proxy $4n\sigma^{2}\|f-g\|_{L_{2}(\mathbb{P}_{n})}^{2}$. A sub-Gaussian random variable $Z$ with mean $\rho<0$ and variance proxy $\tau^{2}$ can be shown to satisfy $P(Z>0)\leq\exp(-\rho^{2}/2\tau^{2})$ using arguments from boucheron2013. Applying this result, as in the previous lemma, we conclude that conditional on $\vec{X}$ and the event $E$, $\psi=1$ holds with probability bounded by
$\exp\left(-\frac{\left(\sqrt{C^{2}-1}-2\sqrt{2}\right)^{2}\delta^{2}}{8\sigma^{2}}\right).$
The rest of the proof follows identically to Lemma 3.2. Then we use this generalized lemma to obtain Lemma 3.3 since normality is not used. ∎
Thus we obtain similar algorithmic guarantees in the following Theorem. By Remark 3.4, the proof holds for sub-Gaussian noise as well.
Theorem 3.5.
Suppose $\mathcal{F}$ satisfies (3.1). Let $J^{\ast}$ be the maximal integer $J\geq 1$ with $\varepsilon_{J}=\frac{d\sqrt{\widetilde{L}(c/2-1,\sigma)}}{2^{J-2}c}$ satisfying
$n\varepsilon_{J}^{2}>2\log M_{\mathcal{F}}^{\operatorname{loc}}\left(\frac{\varepsilon_{J}\cdot c}{\sqrt{\widetilde{L}(c/2-1,\sigma)}},c\right)\vee\log 2.$ (3.5)
and $J^{\ast}=1$ if this condition is never satisfied. Let $f^{\ast}(\vec{X},\vec{Y})$ be the output of $J^{\ast}$ steps of Algorithm 1. Then $\mathbb{E}_{X,Y}\|\bar{f}-f^{\ast}\|_{L_{2}(\mathbb{P}_{X})}^{2}\leq\overline{C}\varepsilon_{J^{\ast}}^{2}$ for some constant $\overline{C}$ depending only on universal constants and $\sigma$.We defined $\widetilde{L}(C,\sigma)$ in (3.3), where $C$ is some sufficiently large universal constant and $c:=2(C+1)$ is the constant we use in the definition of local metric entropy.
{toappendix}
Proof of Theorem 3.5.
Suppose $J>1$ steps of the algorithm are performed and consider $2\leq j\leq J$. Let $\Upsilon_{j+1}$ be the function added to $\Upsilon$ in step $j$. During step $j-1$, we start with some $\Upsilon_{j-1}$ and then form a $\frac{d}{2^{j-1}(C+1)}$-packing $M_{j-1}$ of the set $B(\Upsilon_{j-1},d/2^{j-2})\cap\mathcal{F}$. We then set $\Upsilon_{j}=\mathop{\mathrm{argmin}}_{f\in M_{j-1}}\sum_{i=1}^{n}(Y_{i}-f(X_{i}))^{2}$. Applying Lemma 3.3 with $\delta=\frac{d}{2^{j-1}(C+1)}$, $\mathcal{F^{\prime}}:=B(\Upsilon_{j-1},d/2^{j-2})\cap\mathcal{F}$, and conditioning on $\|\bar{f}-\Upsilon_{j-1}\|_{L_{2}(\mathbb{P}_{X})}\leq d/2^{j-2}$ (which implies $\bar{f}\in\mathcal{F}^{\prime}$), we obtain
$\mathbb{P}_{\bar{f}}\left(\|\bar{f}-\Upsilon_{j}\|_{L_{2}(\mathbb{P}_{X})}\geq\frac{d}{2^{j-1}}\bigg{|}\|\bar{f}-\Upsilon_{j-1}\|_{L_{2}(\mathbb{P}_{X})}\leq\frac{d}{2^{j-2}},\Upsilon_{j-1}\right)\leq 3\cdot|M_{j-1}|\exp(-n\varepsilon_{j}^{2}),$
where we define
$\varepsilon_{j}$ $=\frac{d\sqrt{\widetilde{L}(C,\sigma)}}{2^{j-1}(C+1)}.$
Further, define $a_{j}=\exp(-n\varepsilon_{j}^{2})$. Observe that $|M_{j-1}|$ is upper-bounded by $M_{\mathcal{F}}^{\operatorname{loc}}(d/2^{j-2},2(C+1))$, and by Lemma 2.9 this quantity is bounded by $M_{\mathcal{F}}^{\operatorname{loc}}(d/2^{J-2},2(C+1))$. The probability upper bound does not depend on $\Upsilon_{j-1}$, so we have proven that
$\mathbb{P}_{\bar{f}}\left(\|\bar{f}-\Upsilon_{j}\|_{L_{2}(\mathbb{P}_{X})}\geq\frac{d}{2^{j-1}}\bigg{|}\|\bar{f}-\Upsilon_{j-1}\|_{L_{2}(\mathbb{P}_{X})}\leq\frac{d}{2^{j-2}}\right)$ $\leq 3\cdot M_{\mathcal{F}}^{\operatorname{loc}}(d/2^{J-2},2(C+1))\cdot\exp(-n\varepsilon_{j}^{2}).$
If we set $A_{j}$ to be the event that $\|\bar{f}-\Upsilon_{j}\|_{L_{2}(\mathbb{P}_{X})}\geq\frac{d}{2^{j-1}}$, then we can use the property that $P(A_{J})\leq P(A_{J}|A_{J-1}^{c})+P(A_{j-1}|A_{j-2}^{c})+\dots+P(A_{3}|A_{2}^{c})+P(A_{2})$ to conclude that
$\mathbb{P}_{\bar{f}}\left(\|\bar{f}-\Upsilon_{J}\|_{L_{2}(\mathbb{P}_{X})}\geq\frac{d}{2^{J-1}}\right)$ $\leq 3\cdot M_{\mathcal{F}}^{\operatorname{loc}}(d/2^{J-2},2(C+1))\cdot\sum_{j=2}^{J}a_{j}$
$\leq 3\cdot\mathbbm{1}(J>1)M_{\mathcal{F}}^{\operatorname{loc}}(d/2^{J-2},2(C+1))\frac{a_{J}}{1-a_{J}}.$
The final inequality used Lemma 2.10. Note that $a_{J}<1$ since we assume $C$ is large enough that $\widetilde{L}(C,\sigma)\neq 0$.
Note that $\frac{d}{2^{J-2}}=\frac{2\varepsilon_{J}(C+1)}{\sqrt{\widetilde{L}(C,\sigma)}}.$ Now, if $n\varepsilon_{J}^{2}>2\log M_{\mathcal{F}}^{\operatorname{loc}}\left(\frac{2\cdot\varepsilon_{J}(C+1)}{\sqrt{\widetilde{L}(C,\sigma)}},2(C+1)\right)$ and $a_{J}=\exp\left(-n\varepsilon_{J}^{2}\right)<\frac{1}{2}$, which implies $(1-a_{J})^{-1}<2$, then we have
$M_{\mathcal{F}}^{\operatorname{loc}}(d/2^{J-2},2(C+1))\cdot\frac{a_{J}}{1-a_{J}}$ $=M_{\mathcal{F}}^{\operatorname{loc}}\Bigg{(}\frac{2\varepsilon_{J}(C+1)}{\sqrt{\widetilde{L}(C,\sigma)}},2(C+1)\Bigg{)}\cdot\frac{a_{J}}{1-a_{J}}$
$<\exp\left(\frac{n\varepsilon_{J}^{2}}{2}\right)\cdot\frac{a_{J}}{1-a_{J}}$
$\leq\exp\left(\frac{n\varepsilon_{J}^{2}}{2}\right)\cdot\exp\left(-n\varepsilon_{J}^{2}\right)\cdot\frac{1}{1-a_{J}}$
$\leq 2\exp\left(-\frac{n\varepsilon_{J}^{2}}{2}\right).$
Since $a_{J}<\frac{1}{2}$ iff $n\varepsilon_{J}^{2}>\log 2$, we conclude that if
$n\varepsilon_{J}^{2}>2\log M_{\mathcal{F}}^{\operatorname{loc}}\left(\frac{2\cdot\varepsilon_{J}(C+1)}{\sqrt{\widetilde{L}(C,\sigma)}},2(C+1)\right)\vee\log 2,$
then $M_{\mathcal{F}}^{\operatorname{loc}}(d/2^{J-2},2(C+1))\cdot\frac{a_{J}}{1-a_{J}}\leq 2\exp\left(-\frac{n\varepsilon_{J}^{2}}{2}\right)$.
To summarize, if condition (3.5) holds for $\varepsilon_{J}$, then we have
$\mathbb{P}_{\bar{f}}\left(\|\bar{f}-\Upsilon_{J}\|_{L_{2}(\mathbb{P}_{X})}\geq\frac{d}{2^{J-1}}\right)$ $\leq 3\cdot\mathbbm{1}(J>1)M_{\mathcal{F}}^{\operatorname{loc}}(d/2^{J-2},2(C+1))\frac{a_{J}}{1-a_{J}}$
$\leq 6\cdot\mathbbm{1}(J>1)\exp\left(-\frac{n\varepsilon_{J}^{2}}{2}\right).$
Thus, (3.5) implies a lower bound on $P(A_{J}^{c})$.
Now, set $\kappa=\frac{3(C+1)}{\sqrt{\widetilde{L}(C,\sigma)}}$. Suppose $J$ satisfies condition (3.5). Then $J\leq J^{\ast}$ with $f^{\ast}$ the output of $J^{\ast}$ steps of the algorithm. By the Cauchy sequence argument in (2.2), we have $\|f^{\ast}-\Upsilon_{J}\|_{L_{2}(\mathbb{P}_{X})}\leq\frac{d}{2^{J-2}}$. Now if we assume further that $\|\Upsilon_{K}-\bar{f}\|_{L_{2}(\mathbb{P}_{X})}\leq\frac{d}{2^{J-1}}$, i.e., the event $A_{J}^{c}$, we have by the triangle inequality that $\|f^{\ast}-\bar{f}\|_{L_{2}(\mathbb{P}_{X})}\leq\frac{3d}{2^{J-1}}=\kappa\varepsilon_{J}$. If we call $B_{J}$ the event that $\|f^{\ast}-\bar{f}\|_{L_{2}(\mathbb{P}_{X})}\leq\kappa\varepsilon_{J}$, we have $P(A_{J}^{c})\leq P(B_{J})$ for $J$ satisfying (3.5). We already produced a lower bound on $P(A_{J}^{c})$ for such $J$. So in fact, for all $J$ satisfying (3.5), we have a lower bound on $P(B_{J})$, namely, that
$\|\bar{f}-f^{\ast}\|_{L_{2}(\mathbb{P}_{X})}$ $\leq\kappa\cdot\varepsilon_{J}$
with probability at least
$1-6\cdot\exp\left(-\frac{n\varepsilon_{J}^{2}}{2}\right).$
Let $J^{\ast}$ be the maximal $J$ for which conditions (3.5) all hold, with $J^{\ast}=1$ if none do. Note that $\varepsilon_{J}\downarrow 0$ as $J\to\infty$, so (3.5) will not hold for all sufficiently large $J$. This implies $J^{\ast}<\infty$. Further, set $\underline{C}=6$ and $C^{\prime}=\frac{1}{2}$. Then for any $1\leq J\leq J^{\ast}$ including $J^{\ast}=1$, we have
$\mathbb{P}_{\bar{f}}\left(\|\bar{f}-f^{\ast}\|_{L_{2}(\mathbb{P}_{X})}\geq\kappa\cdot\varepsilon_{J}\right)$ $\leq\underline{C}\cdot\mathbbm{1}(J>1)\exp(-C^{\prime}n\varepsilon_{J}^{2})$
$\leq\underline{C}\cdot\mathbbm{1}(J^{\ast}>1)\exp(-C^{\prime}n\varepsilon_{J}^{2})$ (3.6)
Note the first inequality trivially holds for $J=1$ since $\kappa\varepsilon_{1}=3d$, so that the probability will be 0 as will the right-hand side. Similarly, we extend this to hold for all $J\in\mathbb{Z}$ with $J\leq J^{\ast}$, as if $J\leq 0$, then $\kappa\cdot\varepsilon_{J}=\frac{3d}{2^{J-1}}\geq 6d$. But since $d$ is the diameter of $\mathcal{F}$, the above probability is 00 which is less than than the right hand side of (3.6).
Now observe that $\bigcup_{J\leq J^{\ast}}[\varepsilon_{J},\varepsilon_{J-1})=[\varepsilon_{J^{\ast}},\infty)$. Pick any $x\geq\varepsilon_{J^{\ast}}$. Then $x\in[\varepsilon_{J},\varepsilon_{J-1})$ for some $J\leq J^{\ast}$. Then
$2\kappa x\geq 2\kappa\cdot\varepsilon_{J}=2\kappa\cdot\varepsilon_{J-1}.$
Note further that $x\mapsto\exp(-C^{\prime}nx^{2})$ is non-increasing for $x\geq 0$. Consequently,
$\mathbb{P}_{\bar{f}}\left(\|\bar{f}-f^{\ast}\|_{L_{2}(\mathbb{P}_{X})}\geq 2\kappa x\right)$ $\leq\mathbb{P}_{\bar{f}}\left(\|\bar{f}-f^{\ast}\|_{L_{2}(\mathbb{P}_{X})}\geq\kappa\cdot\varepsilon_{J-1}\right)$
$\leq\underline{C}\cdot\mathbbm{1}(J^{\ast}>1)\exp(-C^{\prime}n\varepsilon_{J-1}^{2})$
$\leq\underline{C}\cdot\mathbbm{1}(J^{\ast}>1)\exp(-C^{\prime}nx^{2}).$
Thus $\mathbb{P}_{\bar{f}}\left(\|\bar{f}-f^{\ast}\|_{L_{2}(\mathbb{P}_{X})}\geq 2\kappa x\right)\leq\underline{C}\cdot\mathbbm{1}(J^{\ast}>1)\exp(-C^{\prime}nx^{2})$ for all $x\geq\varepsilon_{J^{\ast}}$.
Then we integrate:
$\mathbb{E}_{X,Y}\|\bar{f}-f^{\ast}\|_{L_{2}(\mathbb{P}_{X})}^{2}$ $=\int_{0}^{\infty}2x\mathbb{P}_{\bar{f}}\left(\|\bar{f}-f^{\ast}\|_{L_{2}(\mathbb{P}_{X})}\geq x\right)\mathrm{d}x$
$\leq C^{\prime\prime\prime}{\varepsilon_{J^{\ast}}}^{2}+\int_{2\kappa\varepsilon_{J^{\ast}}}^{\infty}2x\cdot\underline{C}\cdot\mathbbm{1}(J^{\ast}>1)\exp(-C^{\prime\prime}nx^{2})\mathrm{d}x$
$=C^{\prime\prime\prime}{\varepsilon_{J^{\ast}}}^{2}+C^{\prime\prime\prime\prime}\cdot\mathbbm{1}(J^{\ast}>1)n^{-1}\exp(-C^{\prime\prime\prime\prime\prime}n{\varepsilon_{J^{\ast}}}^{2}).$
Recall from our assumption (3.5) that either $J^{\ast}=1$ or $n{\varepsilon_{J^{\ast}}}^{2}>\log 2$. In either case, $\mathbb{E}_{X,Y}\|\bar{f}-f^{\ast}\|_{L_{2}(\mathbb{P}_{X})}^{2}\leq\overline{C}{\varepsilon_{J^{\ast}}}^{2}$ for some constant $\overline{C}$ depending only on $\sigma$, $d$, and our constant $C>3$.
The following theorem then brings together the lower and upper bounds we obtained.
Theorem 3.6.
Suppose $\mathcal{F}$ satisfies (3.1). Define $\varepsilon^{\ast}:=\sup\{\varepsilon:n\varepsilon^{2}\leq\log M_{\mathcal{F}}^{\operatorname{loc}}(\varepsilon,c)\}$ where $c$ is a sufficiently large absolute constant. Then the minimax rate is ${\varepsilon^{\ast}}^{2}\wedge d^{2}$ up to absolute constants.
{toappendix}
Proof of Theorem 3.6.
The proof is entirely identical to Theorem 2.12, except we replace $L(c/2-1,\sigma,B_{\mathcal{F}})$ from the bounded case with $\widetilde{L}(c/2-1,\sigma)$ everywhere. E.g., we set $\eta=\frac{c\sqrt{2}}{\widetilde{L}(c/2-1,\sigma)}\wedge 1\in(0,1)$. ∎
4 Adaptive Version
We relax the definition of local metric entropy by removing the supremum and instead considering the pointwise value of the local metric entropies. This lets us produce an estimator that is adaptive to the true point $\bar{f}$, which is a difference from the work in yang1999information and mendelson_2017_local_versus_global, and to our knowledge is the first example of an adaptive estimator in this general setting.
Definition 4.1 (Adaptive Local Entropy).
Let $M(r,\mathcal{F})$ be the largest cardinality of a $r$-packing set in the $L_{2}(\mathbb{P}_{X})$ norm of $\mathcal{F}$. Let $c>0$. Then we define
$M_{\mathcal{F}}^{\operatorname{adloc}}(f,\varepsilon,c)=M(\varepsilon/c,B(f,\varepsilon)\cap\mathcal{F}).$
Then $\log M^{\operatorname{adloc}}_{\mathcal{F}}(f,\varepsilon,c)$ is the adaptive local entropy of $\mathcal{F}$ at $f$.
A following lemma relates the adaptive local metric entropy of two nearby points.
Lemma 4.2.
Let $f,g\in\mathcal{F}$ satisfy $\|f-g\|_{L_{2}(\mathbb{P}_{X})}<\delta$. Then $M_{\mathcal{F}}^{\operatorname{adloc}}(f,\varepsilon,c)\leq M_{\mathcal{F}}^{\operatorname{adloc}}(g,2\varepsilon,2c)$ provided $\varepsilon>\delta$.
Proof.
If we show $B(f,\varepsilon)\subseteq B(g,2\varepsilon)$, then the cardinality of the largest $(\varepsilon/c)$-packing of $B(f,\varepsilon)\cap\mathcal{F}$ will be less than or equal to the cardinality of the largest $(2\varepsilon/2c)$-packing of $B(g,2\varepsilon)\cap\mathcal{F}$. Well, if $h\in B(f,\varepsilon)$, then $\|h-f\|_{L_{2}(\mathbb{P}_{X})}<\varepsilon$. Using the triangle inequality,
$\|h-g\|_{L_{2}(\mathbb{P}_{X})}\leq\|h-f\|_{L_{2}(\mathbb{P}_{X})}+\|g-h\|_{L_{2}(\mathbb{P}_{X})}\leq\varepsilon+\delta<2\varepsilon,$
showing that $h\in B(g,2\varepsilon)$. ∎
We now show that the main estimator is adaptive to the true point $\bar{f}$, in which the key condition (4.2) does not require taking a supremum over all $f\in\mathcal{F}$ in the metric entropy term. The proof works no matter if $\mathcal{F}$ is bounded or not in sup-norm and also applies in the sub-Gaussian noise setting.
Theorem 4.3.
Consider the bounded case. Let $\overline{J}$ be the maximal integer $J\geq 1$ with $\varepsilon_{J}=\frac{d\sqrt{L(c/2-1,\sigma,B_{\mathcal{F}})}}{2^{J-2}c}$ satisfying
$n\varepsilon_{J}^{2}>\inf_{f\in\mathcal{F}}2\log M_{\mathcal{F}}^{\operatorname{adloc}}\left(f,\frac{2\varepsilon_{J}\cdot c}{\sqrt{L(c/2-1,\sigma,B_{\mathcal{F}})}},2c\right)\vee\log 2.$ (4.1)
and $\overline{J}=1$ if this condition is never satisfied. Let $f^{\ast}(\vec{X},\vec{Y})$ be the output of $\overline{J}$ steps of Algorithm 1.
Next, let $J^{\ast}$ be the maximal integer $J\geq 1$ with $\varepsilon_{J}$ satisfying
$n\varepsilon_{J}^{2}>2\log M_{\mathcal{F}}^{\operatorname{adloc}}\left(\bar{f},\frac{2\varepsilon_{J}\cdot c}{\sqrt{L(c/2-1,\sigma,B_{\mathcal{F}})}},2c\right)\vee\log 2,$ (4.2)
noting the use of the true $\bar{f}\in\mathcal{F}$. Then $\mathbb{E}_{X,Y}\|\bar{f}-f^{\ast}\|_{L_{2}(\mathbb{P}_{X})}^{2}\leq\overline{C}{\varepsilon_{J^{\ast}}}^{2}$ for some constant $\overline{C}$ depending only on universal constants, $B_{\mathcal{F}}$, and $\sigma$. We defined $L(C,\sigma,B_{\mathcal{F}})$ in (2.3), where $C$ is some sufficiently large universal constant and $c:=2(C+1)$ is the constant we use in the definition of local metric entropy.
In the unbounded case, assuming $\mathcal{F}$ satisfies (3.1), simply replace every instance of $L(C,\sigma,B_{\mathcal{F}})$ with $\widetilde{L}(C,\sigma)$, which is defined in (3.3).
{toappendix}
Proof of Theorem 4.3.
Suppose $J>1$ steps of the algorithm are performed and consider $2\leq j\leq J$. Let $\Upsilon_{j+1}$ be the function added to $\Upsilon$ in step $j$. During step $j-1$, we start with some $\Upsilon_{j-1}$ and then form a $\frac{d}{2^{j-1}(C+1)}$-packing $M_{j-1}$ of the set $B(\Upsilon_{j-1},d/2^{j-2})\cap\mathcal{F}$. We then set $\Upsilon_{j}=\mathop{\mathrm{argmin}}_{f\in M_{j-1}}\sum_{i=1}^{n}(Y_{i}-f(X_{i}))^{2}$.
Applying Lemma 2.7 with $\delta=\frac{d}{2^{j-1}(C+1)}$, $\mathcal{F^{\prime}}:=B(\Upsilon_{j-1},d/2^{j-2})\cap\mathcal{F}$, and conditioning on $\|\bar{f}-\Upsilon_{j-1}\|_{L_{2}(\mathbb{P}_{X})}\leq d/2^{j-2}$ (which implies $\bar{f}\in\mathcal{F}^{\prime}$), we obtain
$\mathbb{P}_{\bar{f}}\left(\|\bar{f}-\Upsilon_{j}\|_{L_{2}(\mathbb{P}_{X})}\geq\frac{d}{2^{j-1}}\bigg{|}\|\bar{f}-\Upsilon_{j-1}\|_{L_{2}(\mathbb{P}_{X})}\leq\frac{d}{2^{j-2}},\Upsilon_{j-1}\right)$ $\leq 3\cdot|M_{j-1}|\exp(-n\varepsilon_{j}^{2})$
where we define
$\varepsilon_{j}$ $=\frac{d\sqrt{L(C,\sigma,B_{\mathcal{F}})}}{2^{j-1}(C+1)}.$
Further, define $a_{j}=\exp(-n\varepsilon_{j}^{2})$.
Now, we upper bound $|M_{j-1}|$ using Lemma 4.2. In the notation of the lemma, we would set $\varepsilon=\frac{d}{2^{j-2}}$, $\delta=\frac{d}{2^{j-1}}$, $f=\Upsilon_{j-1}$, and $g=\bar{f}$. Then $M_{j-1}$ corresponds to a $\varepsilon/c$-packing of $B(f,\varepsilon)$, and we further condition on $\|\bar{f}-\Upsilon_{j-1}\|_{L_{2}(\mathbb{P}_{X})}=\|g-f\|_{L_{2}(\mathbb{P}_{X})}\leq\delta$, noting that $\delta<\varepsilon$. Thus Lemma 4.2 implies
$|M_{j-1}|$ $\leq M_{\mathcal{F}}^{\operatorname{adloc}}\left(\bar{f},d/2^{j-3},4(C+1)\right)$
$\leq M_{\mathcal{F}}^{\operatorname{adloc}}\left(\bar{f},d/2^{J-3},4(C+1)\right).$
The second inequality uses Lemma 2.9 since $j\leq J$. The resulting probability upper bound does not depend on $\Upsilon_{j-1}$, so we have proven that
$\mathbb{P}_{\bar{f}}\left(\|\bar{f}-\Upsilon_{j}\|_{L_{2}(\mathbb{P}_{X})}\geq\frac{d}{2^{j-1}}\bigg{|}\|\bar{f}-\Upsilon_{j-1}\|_{L_{2}(\mathbb{P}_{X})}\leq\frac{d}{2^{j-2}}\right)$ $\leq 3\cdot M_{\mathcal{F}}^{\operatorname{adloc}}\left(\bar{f},d/2^{J-3},4(C+1)\right)\exp(-n\varepsilon_{j}^{2}).$
If we set $A_{j}$ to be the event $\|\bar{f}-\Upsilon_{j}\|_{L_{2}(\mathbb{P}_{X})}\geq\frac{d}{2^{j-1}}$, then we can use the property that $P(A_{J})\leq P(A_{J}|A_{J-1}^{c})+P(A_{j-1}|A_{j-2}^{c})+\dots+P(A_{3}|A_{2}^{c})+P(A_{2})$ to conclude that
$\mathbb{P}_{\bar{f}}\left(\|\bar{f}-\Upsilon_{J}\|_{L_{2}(\mathbb{P}_{X})}\geq\frac{d}{2^{J-1}}\right)$ $\leq 3\cdot M_{\mathcal{F}}^{\operatorname{adloc}}\left(\bar{f},d/2^{J-3},4(C+1)\right)\cdot\sum_{j=2}^{J}a_{j}$
$\leq 3\cdot\mathbbm{1}(J>1)M_{\mathcal{F}}^{\operatorname{adloc}}\left(\bar{f},d/2^{J-3},4(C+1)\right)\frac{a_{J}}{1-a_{J}}.$
The final inequality used Lemma 2.10. Note that $a_{J}<1$ since $C$ is large enough that $L(C,\sigma,B_{\mathcal{F}})\neq 0$.
Note that
$\frac{d}{2^{J-3}}=\frac{4\varepsilon_{J}(C+1)}{\sqrt{L(C,\sigma,B_{\mathcal{F}})}}.$
Now, if $n\varepsilon_{J}^{2}>2\log M_{\mathcal{F}}^{\operatorname{adloc}}\left(\bar{f},\frac{4\cdot\varepsilon_{J}(C+1)}{\sqrt{L(C,\sigma,B_{\mathcal{F}})}},4(C+1)\right)$ and $a_{J}=\exp\left(-n\varepsilon_{J}^{2}\right)<\frac{1}{2}$, which implies $(1-a_{J})^{-1}<2$, then we have
$M_{\mathcal{F}}^{\operatorname{adloc}}\left(\bar{f},\frac{d}{2^{J-3}},4(C+1)\right)\cdot\frac{a_{J}}{1-a_{J}}$ $=M_{\mathcal{F}}^{\operatorname{adloc}}\left(\bar{f},\frac{4\varepsilon_{J}(C+1)}{\sqrt{L(C,\sigma,B_{\mathcal{F}})}},4(C+1)\right)\cdot\frac{a_{J}}{1-a_{J}}$
$<\exp\left(\frac{n\varepsilon_{J}^{2}}{2}\right)\cdot\frac{a_{J}}{1-a_{J}}$
$\leq\exp\left(\frac{n\varepsilon_{J}^{2}}{2}\right)\cdot\exp\left(-n\varepsilon_{J}^{2}\right)\cdot\frac{1}{1-a_{J}}$
$\leq 2\exp\left(-\frac{n\varepsilon_{J}^{2}}{2}\right).$
Since $a_{J}<\frac{1}{2}$ iff $n\varepsilon_{J}^{2}>\log 2$, we conclude that if
$n\varepsilon_{J}^{2}>2\log M_{\mathcal{F}}^{\operatorname{adloc}}\left(\bar{f},\frac{4\cdot\varepsilon_{J}(C+1)}{\sqrt{L(C,\sigma,B_{\mathcal{F}})}},4(C+1)\right)\vee\log 2,$
then $M_{\mathcal{F}}^{\operatorname{adloc}}(\bar{f},d/2^{J-3},4(C+1))\cdot\frac{a_{J}}{1-a_{J}}\leq 2\exp\left(-\frac{n\varepsilon_{J}^{2}}{2}\right)$.
To summarize, if condition (4.2) holds for $\varepsilon_{J}$, then we have the following upper bound on $P(A_{J})$:
$\mathbb{P}_{\bar{f}}\left(\|\bar{f}-\Upsilon_{J}\|_{L_{2}(\mathbb{P}_{X})}\geq\frac{d}{2^{J-1}}\right)$ $\leq 3\cdot\mathbbm{1}(J>1)M_{\mathcal{F}}^{\operatorname{adloc}}\left(\bar{f},\frac{d}{2^{J-3}},4(C+1)\right)\frac{a_{J}}{1-a_{J}}$
$\leq 6\cdot\mathbbm{1}(J>1)\exp\left(-\frac{n\varepsilon_{J}^{2}}{2}\right).$
This equivalently gives us a lower bound on $P(A_{J}^{c})$ of $1-6\cdot\mathbbm{1}(J>1)\exp\left(-\frac{n\varepsilon_{J}^{2}}{2}\right)$.
Now, set $\kappa=\frac{3(C+1)}{\sqrt{L(C,\sigma,B_{\mathcal{F}})}}$. Suppose $J$ satisfies (4.2). Then $J\leq J^{\ast}$ and by definition $J^{\ast}\leq\overline{J}$. Recall also that $f^{\ast}$ was the output after $\overline{J}$ steps. Hence by the Cauchy sequence argument in (2.2), since $J\leq\overline{J}$, we have $\|f^{\ast}-\Upsilon_{J}\|_{L_{2}(\mathbb{P}_{X})}\leq\frac{d}{2^{J-2}}$. If we assume further that $\|\Upsilon_{J}-\bar{f}\|_{L_{2}(\mathbb{P}_{X})}\leq\frac{d}{2^{J-1}}$, then by the triangle inequality, $\|\bar{f}-f^{\ast}\|_{L_{2}(\mathbb{P}_{X})}\leq\frac{3d}{2^{J-1}}=\kappa\cdot\varepsilon_{J}$. Now call $B_{J}$ the event that $\|f^{\ast}-\bar{f}\|_{L_{2}(\mathbb{P}_{X})}\leq\kappa\varepsilon_{J}$, and notice we have $P(A_{J}^{c})\leq P(B_{J})$ for such $J$ satisfying (4.2). We already gave a lower bound on $P(A_{J}^{c})$ for such $J$. Thus, we have proven for all $J$ satisfying (4.2) that
$\|\bar{f}-f^{\ast}\|_{L_{2}(\mathbb{P}_{X})}$ $\leq\kappa\cdot\varepsilon_{J}$
holds with probability at least
$1-6\cdot\exp\left(-\frac{n\varepsilon_{J}^{2}}{2}\right).$
Let $J^{\ast}$ be the maximal $J$ for which conditions (4.2) all hold, with $J^{\ast}=1$ if none do. Note that $\varepsilon_{J}\downarrow 0$ as $J\to\infty$, so (4.2) will not hold for all sufficiently large $J$. This implies $J^{\ast}<\infty$. Further, set $\underline{C}=6$, $C^{\prime}=\frac{1}{2}$. Then for any $1\leq J\leq J^{\ast}$ including $J^{\ast}=1$, we have
$\mathbb{P}_{\bar{f}}\left(\|\bar{f}-f^{\ast}\|_{L_{2}(\mathbb{P}_{X})}\geq\kappa\cdot\varepsilon_{J}\right)$ $\leq\underline{C}\cdot\mathbbm{1}(J>1)\exp(-C^{\prime}n\varepsilon_{J}^{2})$
$\leq\underline{C}\cdot\mathbbm{1}(J^{\ast}>1)\exp(-C^{\prime}n\varepsilon_{J}^{2})$ (4.3)
The first inequality holds trivially for $J=1$ as $\kappa\varepsilon_{1}=3d$ and $d$ is the diameter of $\mathcal{F}$, so both sides of the inequality are 0. Similarly we extend this result to all $J\in\mathbb{Z}$ with $J\leq J^{\ast}$, as if $J\leq 0$, then $\kappa\cdot\varepsilon_{J}=\frac{3d}{2^{J-1}}\geq 6d$ so the left-hand side of the first inequality is 0.
Next, suppose $x\in[\varepsilon_{J},\varepsilon_{J-1})$. Then
$2\kappa x\geq 2\kappa\cdot\varepsilon_{J}=2\kappa\cdot\varepsilon_{J-1}.$
Note further that $x\mapsto\exp(-C^{\prime}nx^{2})$ is non-increasing for $x\geq 0$. Consequently,
$\mathbb{P}_{\bar{f}}\left(\|\bar{f}-f^{\ast}\|_{L_{2}(\mathbb{P}_{X})}\geq 2\kappa x\right)$ $\leq\mathbb{P}_{\bar{f}}\left(\|\bar{f}-f^{\ast}\|_{L_{2}(\mathbb{P}_{X})}\geq\kappa\cdot\varepsilon_{J-1}\right)$
$\leq\underline{C}\cdot\mathbbm{1}(J^{\ast}>1)\exp(-C^{\prime}n\varepsilon_{J-1}^{2})$
$\leq\underline{C}\cdot\mathbbm{1}(J^{\ast}>1)\exp(-C^{\prime}nx^{2}).$
Then integrate:
$\mathbb{E}_{X,Y}\|\bar{f}-f^{\ast}\|_{L_{2}(\mathbb{P}_{X})}^{2}$ $=\int_{0}^{\infty}2x\mathbb{P}_{\bar{f}}\left(\|\bar{f}-f^{\ast}\|_{L_{2}(\mathbb{P}_{X})}\geq x\right)\mathrm{d}x$
$\leq C^{\prime\prime\prime}{\varepsilon_{J^{\ast}}}^{2}+\int_{2\kappa\varepsilon_{J^{\ast}}}^{\infty}2x\cdot\underline{C}\cdot\mathbbm{1}(J^{\ast}>1)\exp(-C^{\prime\prime}nx^{2})\mathrm{d}x$
$=C^{\prime\prime\prime}{\varepsilon_{J^{\ast}}}^{2}+C^{\prime\prime\prime\prime}\cdot\mathbbm{1}(J^{\ast}>1)n^{-1}\exp(-C^{\prime\prime\prime\prime\prime}n{\varepsilon_{J^{\ast}}}^{2}).$
Recall from our assumption (4.2) that either $J^{\ast}=1$ or $n{\varepsilon_{J^{\ast}}}^{2}>\log 2$. In either case, $\mathbb{E}_{X,Y}\|\bar{f}-f^{\ast}\|_{L_{2}(\mathbb{P}_{X})}^{2}\leq\overline{C}{\varepsilon_{J^{\ast}}}^{2}$ for some constant $\overline{C}$ depending only on $\sigma$, $B_{\mathcal{F}}$, and our constant $C>3$.
5 Examples
We start out by considering function spaces defined as linear functionals over compact, convex sets $K$ using sub-Gaussian isotropic data; such function spaces will not have a uniform bound on the sup-norm. For such spaces, local metric entropies in our function space using an $L_{2}(\mathbb{P}_{X})$-norm will be equivalent to local metric entropies in $\mathbb{R}^{p}$ with the Euclidean 2-norm. We then specialize our results to the $\ell_{1}$-ball and ellipsoids. We later consider functions with a smoothness condition as well as consider monotone functions in a multivariate setting.
5.1 Linear Functionals
We call a random variable $Z\in\mathbb{R}^{p}$ isotropic if $\mathbb{E}[Z]=0,\mathbb{E}[ZZ^{T}]=\mathbf{I}_{p}$. Moreover, we say $Z\in\mathbb{R}^{p}$ is a sub-Gaussian random vector with mean $\mu\in\mathbb{R}^{p}$ and variance proxy $\tau^{2}$ if for any $v\in\mathbb{R}^{p}$ with $\|v\|_{2}=1$ we have $v^{T}(Z-\mu)$ is a sub-Gaussian random variable with mean 00 and variance proxy $\tau^{2}$.
Let $K\subset\mathbb{R}^{p}$ be a compact, convex set. Define =$\mathcal{F}=\{f_{\beta}\colon\mathbb{R}^{p}\to\mathbb{R},f_{\beta}(x)=x^{T}\beta:\beta\in K\}$. Now let us draw i.i.d. isotopric sub-Gaussian vectors $X_{1},\dots,X_{n}$ with mean 00 and variance proxy $\tau^{2}>1$. Importantly, functions in $\mathcal{F}$ are not bounded in sup-norm despite $K$ being compact. Observe that
$\mathrm{diam}(\mathcal{F})^{2}$ $=\sup_{f_{\beta_{1}},f_{\beta_{2}}\in\mathcal{F}}\|f_{\beta_{1}}-f_{\beta_{2}}\|_{L_{2}(\mathbb{P}_{X})}^{2}=\sup_{\beta_{1},\beta_{2}\in K}\mathbb{E}\left[(\beta_{1}-\beta_{2})^{T}X\right]^{2}$
$=\sup_{\beta_{1},\beta_{2}\in K}\|\beta_{1}-\beta_{2}\|_{2}^{2}=\mathrm{diam}(K)^{2}.$
The third equality used the fact that $X$ is isotropic. Thus, the diameter of $\mathcal{F}$ in the $L_{2}(\mathbb{P}_{X})$-norm is the diameter of $K$ which is finite.
Next, note that for any distinct $\beta_{1},\beta_{2}\in K$, we have $\|\beta_{1}-\beta_{2}\|_{2}^{-1}\cdot(\beta_{1}-\beta_{2})^{T}X\sim\mathrm{SG}(0,\tau^{2})$. Let $W$ denote this quantity. By vershynin2018high, for some absolute constant $\alpha>0$ we have $\mathbb{E}|W^{p}|\leq\tau^{p}\alpha^{p}p^{p/2}$. Then we compute that
$\|f_{\beta_{1}}-f_{\beta_{2}}\|_{L_{p}(\mathbb{P}_{X})}^{p}$ $=\mathbb{E}\left[(\beta_{1}-\beta_{2})^{T}X\right]^{p}=\|\beta_{1}-\beta_{2}\|_{2}^{p}\cdot\mathbb{E}[W^{p}]\leq\|\beta_{1}-\beta_{2}\|_{2}^{p}\cdot\mathbb{E}|W^{p}|$
$\leq\|\beta_{1}-\beta_{2}\|_{2}^{p}\cdot\tau^{p}\alpha^{p}p^{p/2}.$
Taking the $p$th root and noting by our above argument that $X$ being isotropic implies $\|\beta_{1}-\beta_{2}\|_{2}=\|f_{\beta_{1}}-f_{\beta_{2}}\|_{L_{2}(\mathbb{P}_{X})}$, we have
$\|f_{\beta_{1}}-f_{\beta_{2}}\|_{L_{p}(\mathbb{P}_{X})}^{p}\leq\tau\alpha\|f_{\beta_{1}}-f_{\beta_{2}}\|_{L_{2}(\mathbb{P}_{X})}\sqrt{p}.$
Hence the condition in (3.1) holds.
Because $\|f_{\beta_{1}}-f_{\beta_{2}}\|_{L_{2}(\mathbb{P}_{X})}=\|\beta_{1}-\beta_{2}\|_{2}$ holds for all $\beta_{1},\beta_{2}\in K$ and since $B(f_{\beta},\varepsilon,\|\cdot\|_{L_{2}(\mathbb{P}_{X})})=B(\beta,\varepsilon,\|\cdot\|_{2})$, it follows that any $\varepsilon/c$-packing of $B(f_{\beta},\varepsilon,\|\cdot\|_{L_{2}(\mathbb{P}_{X})})\cap\mathcal{F}$ corresponds to a $\varepsilon/c$-packing of $B(\beta,\varepsilon,\|\cdot\|_{2})\cap K$ in the Euclidean 2-norm on $\mathbb{R}^{p}$. Hence $M_{\mathcal{F}}^{\operatorname{loc}}(\varepsilon,c)=M_{K}^{\operatorname{loc}}(\varepsilon,c)$.
5.1.1 Linear Functionals on the $\ell_{1}$-ball
Let $K=B_{1}^{p}(0,1):=\{x\in\mathbb{R}^{p}:\|x\|_{1}\leq 1\}$ be the $\ell_{1}$-unit-ball, and let $B_{2}^{p}(x,r)=\{y\in\mathbb{R}^{p}:\|x-y\|_{2}\leq r\}$. Let $\mathbb{S}_{1}^{d-1}=\{x\in\mathbb{R}^{p}:\|x\|_{1}=1\}$ be the $\ell_{1}$-unit sphere, and similarly set $\mathbb{S}_{2}^{p-1}=\{x\in\mathbb{R}^{p}:\|x\|_{2}=1\}$. We again take $\mathcal{F}$ to be the set of linear functionals over $K$. Note that the equivalence of local metric entropies between our function space and $K$ in our linear functionals setting with isotropic and sub-Gaussian data also applies to the adaptive local metric entropy. Suppose $\beta^{\ast}$ is $s$-sparse, i.e., has at most $s$ non-zero components, and that $\beta^{\ast}\in\mathbb{S}_{1}^{p-1}$. Our goal will be to use the adaptive local entropy to derive a minimax rate upper bound for $\mathcal{F}$ under this sparsity assumption. Our result will agree with the work of raskutti2011minimax.
We will need some tools from chandrasekaran2012convex. The authors begin with an atomic set $\mathcal{A}$ which is a compact subset of $\mathbb{R}^{p}$. The authors then construct an atomic norm induced by this atomic set using the gauge function, assuming $\mathcal{A}$ is centrally symmetric about the origin. For our example, we take $\mathcal{A}$ to be the set of $1$-sparse vectors, so that the convex hull $\mathrm{conv}(\mathcal{A})$ is precisely the $\ell_{1}$-unit-ball $B_{1}^{p}(0,1)$. We then define the atomic norm induced by $\mathcal{A}$ at a vector $x$ by $\|x\|_{\mathcal{A}}=\inf\{t>0:x\in t\cdot\mathrm{conv}(\mathcal{A})\}$. Then clearly $\|x\|_{\mathcal{A}}=\|x\|_{1}$ for $x\neq 0$, since the smallest $t>0$ such that $x/t\in B_{1}^{p}(0,1)$ is $\|x\|_{1}$, and $\|0\|_{\mathcal{A}}=\|0\|_{1}$ since $0\in t\cdot B_{1}^{p}(0,1)$ for arbitrarily small $t>0$. Thus, our atomic set and norm exactly correspond to the $\ell_{1}$-ball and $\ell_{1}$-norm.
Next, recalling that the cone of a set $S\subseteq\mathbb{R}^{p}$ is defined by $\mathrm{cone}(S):=\alpha S:=\{\alpha\cdot s:s\in S,\alpha>0\}$, we define the tangent cone at $x\in\mathbb{R}^{p}$ as
$T_{\mathcal{A}}(x)=\mathrm{cone}\{z-x:\|z\|_{\mathcal{A}}\leq\|x\|_{\mathcal{A}}\}=\mathrm{cone}\{z-x:\|z\|_{1}\leq\|x\|_{1}\}.$
Let us compare the set $S_{1}:=\frac{B_{2}^{p}(\beta^{\ast},\varepsilon)\cap B_{1}^{p}(0,1)-\beta^{\ast}}{\varepsilon}$ to $S_{2}:=T_{\mathcal{A}}(\beta^{\ast})\cap B_{2}^{p}(0,1)$, and show $S_{1}\subseteq S_{2}$. Suppose $x\in S_{1}$. Then $\varepsilon x+\beta^{\ast}\in B_{2}^{p}(\beta^{\ast},\varepsilon)\cap B_{1}^{p}(0,1)$. This implies that $\|\varepsilon x+\beta^{\ast}\|_{1}\leq 1=\|\beta^{\ast}\|_{1}$ and $\|\varepsilon x\|_{2}\leq\varepsilon$. The latter implies $\|x\|_{2}\leq 1$, i.e., $x\in B_{2}^{p}(0,1)$. Next observe that
$x=\frac{1}{\varepsilon}\left(\varepsilon x+\beta^{\ast}-\beta^{\ast}\right)\in\mathrm{cone}\{z-\beta^{\ast}:\|z\|_{1}\leq\|\beta^{\ast}\|_{1}\}=T_{\mathcal{A}}(\beta^{\ast}),$
where we choose $z=\varepsilon x+\beta^{\ast}$. Thus, we conclude $x\in S_{2}$.
For any set $T$, define its Gaussian width $w(T)$ by $w(T):=\mathbb{E}_{g\sim N(0,\mathbf{I}_{p})}\left[\sup_{t\in T}\langle t,g\rangle\right].$ We will use numerous properties of the Gaussian width. Namely, $w$ is invariant under translations, is positive homogeneous, i.e., $w(\alpha T)=\alpha w(T)$ for any $\alpha>0$, and if $T_{1}\subseteq T_{2}$, then $w(T_{1})\leq w(T_{2})$. Since the family of random variables $\{\langle t,g\rangle:t\in T\}$ is an example of a zero-mean Gaussian process, we apply the Sudakov minoration lower bound (wainwright2019high, Theorem 5.30) to upper bound the local metric entropy.
$\log M_{K}^{\operatorname{adloc}}(\beta^{\ast},\varepsilon,c)$ $=\log M(\varepsilon/c,B_{2}^{p}(\beta^{\ast},\varepsilon)\cap B_{1}^{p}(0,1))$
$\leq\frac{2c^{2}}{\varepsilon^{2}}\cdot\left[w(B_{2}^{p}(\beta^{\ast},\varepsilon)\cap B_{1}^{p}(0,1))\right]^{2}$
$=\frac{2c^{2}}{\varepsilon^{2}}\cdot\left[w(B_{2}^{p}(\beta^{\ast},\varepsilon)\cap B_{1}^{p}(0,1)-\beta^{\ast})\right]^{2}$
$=2c^{2}\cdot w^{2}\left(\frac{B_{2}^{p}(\beta^{\ast},\varepsilon)\cap B_{1}^{p}(0,1)-\beta^{\ast}}{\varepsilon}\right)$
$\leq 2c^{2}\left[w(T_{\mathcal{A}}(\beta^{\ast})\cap B_{2}^{p}(0,1))\right]^{2}.$
We now relate the Gaussian width of $T_{\mathcal{A}}(\beta^{\ast})\cap B_{2}^{p}(0,1)$ to $T_{\mathcal{A}}(\beta^{\ast})\cap\mathbb{S}_{2}^{p}(0,1)$, and do so by using the statistical dimension of a convex cone $C$, as is defined in amelunxen2014living:
$\delta(C)=\mathbb{E}_{g\sim N(0,\mathbf{I}_{p})}\left[\left(\sup_{t\in C\cap B^{p}(0,1)}\langle t,g\rangle\right)^{2}\right].$
amelunxen2014living implies
$\delta(T_{\mathcal{A}}(\beta^{\ast}))$ $\leq[w(T_{\mathcal{A}}(\beta^{\ast})\cap\mathbb{S}_{2}^{p}(0,1))]^{2}+1.$
Additionally by chandrasekaran2012convex,
$[w(T_{\mathcal{A}}(\beta^{\ast})\cap\mathbb{S}_{2}^{p}(0,1))]^{2}$ $\leq 2s\log\left(\frac{p}{s}\right)+\frac{5}{4}s,$
where we recall $\beta^{\ast}$ is $s$-sparse. Using Jensen’s inequality and the fact that $T_{\mathcal{A}}(\beta^{\ast})$ is a convex cone along with the previous two results, we obtain
$\log M_{K}^{\operatorname{adloc}}(\beta^{\ast},\varepsilon,c)$ $\leq 2c^{2}\left[w(T_{\mathcal{A}}(\beta^{\ast})\cap B_{2}^{p}(0,1))\right]^{2}$
$=2c^{2}\left(\mathbb{E}_{g\sim N(0,\mathbf{I}_{p})}\left[\sup_{t\in T_{\mathcal{A}}(\beta^{\ast})\cap B_{2}^{p}(0,1)}\langle t,g\rangle\right]\right)^{2}$
$\leq 2c^{2}\mathbb{E}_{g\sim N(0,\mathbf{I}_{p})}\left[\left(\sup_{t\in T_{\mathcal{A}}(\beta^{\ast})\cap B_{2}^{p}(0,1)}\langle t,g\rangle\right)^{2}\right]$
$=2c^{2}\delta(T_{\mathcal{A}}(\beta^{\ast})).$
$\leq c^{2}[w(T_{\mathcal{A}}(\beta^{\ast})\cap\mathbb{S}_{2}^{p}(0,1))]^{2}+c^{2}$
$\leq 2c^{2}s\left(2\log\left(\frac{p}{s}\right)+\frac{5}{4}\right)+c^{2}$
$\lesssim s\log(p/s).$
Hence, $\log M_{\mathcal{F}}^{\operatorname{adloc}}(f_{\beta^{\ast}},\varepsilon,c)\lesssim s\log(p/s).$
Now recalling the definition of $\varepsilon_{J}$ and $J^{\ast}$ from Theorem 4.3, we will have for some constant $\kappa>0$ that
$n(\varepsilon_{J^{\ast}}/2)^{2}$ $\leq 2\log M_{\mathcal{F}}^{\operatorname{adloc}}(f_{\beta^{\ast}},\kappa\cdot\varepsilon_{J^{\ast}},c)\vee\log 2\lesssim s\log(p/s).$
Then using Theorem 4.3 and rearranging the previous inequality, we conclude
$\mathbb{E}_{X,Y}\|f_{\beta^{\ast}}-f^{\ast}\|_{L_{2}(\mathbb{P}_{X})}^{2}\lesssim\varepsilon_{J^{\ast}}^{2}\lesssim\frac{s\log(p/s)}{n},$
were $f^{\ast}$ is the output of our algorithm. This result matches that of raskutti2011minimax, although notably our estimator is different, our bound is in expectation rather than in high probability and we use data that are isotropic and sub-Gaussian.
5.1.2 Linear Functionals over Ellipsoids
Define the matrix $K_{c}$ to be the $p\times p$ diagonal matrix with $i$th entry $a_{i}^{-1}$, where $0<a_{1}\leq a_{2}\leq\dots\leq a_{p}$, where $a_{p}\in\mathbb{R}$ is assumed to be a bounded constant independent of the sample size. Let $\Theta=\{\theta\in\mathbb{R}^{p}:\theta^{T}K_{c}\theta\leq 1\}$. Our assumption on $a_{p}$ implies that the set $\Theta$ has a bounded diameter. We draw $X_{1},\dots,X_{n}$ isotropic and sub-Gaussian with variance proxy $\tau^{2}>1$, and for $1\leq i\leq n$ we observe $Y_{i}=X_{i}^{T}\theta^{\ast}+w_{i}$ where $w_{i}\sim N(0,\sigma^{2})$ and $\theta^{\ast}\in\Theta$ is the true value of the parameter. Let $X\in\mathbb{R}^{n\times p}$ be the matrix with $X_{i}$ in the $i$th-row. Our function class is therefore =$\mathcal{F}=\{f_{\theta}(x)=x^{T}\theta:\theta\in\Theta\}$. Note that $\mathrm{diam}(\mathcal{F})=\mathrm{diam}(\Theta)=2\sqrt{a_{p}}$. As we showed in Section 5.1, since the $X_{i}$ are isotropic, $\|f_{\theta_{1}}-f_{\theta_{2}}\|_{L_{2}(\mathbb{P}_{X})}^{2}=\|\theta_{1}-\theta_{2}\|_{2}^{2}$.
Let us now compute the minimax rate for estimating $\theta^{\ast}$, or equivalently, $f_{\theta^{\ast}}$. We will follow neykov2022minimax which has a very similar proof.
Consider the Kolmogorov width $d_{k}(\Theta)=\min_{P\in\mathcal{P}_{k}}\max_{\theta\in\Theta}\|P\theta-\theta\|,$ for $\mathcal{P}_{k}$ the set of linear projections into $k$-dimensional subspaces. Note that $d_{k}(\Theta)=\sqrt{a_{p-k}}$ where we take $a_{0}=0$.
If $a_{p}>1/n$, then let $k\in[p]$ be such that $a_{p-k}\leq(k+1)/n$ and $a_{p-k+1}>k/n$. This will imply $d_{k}(\Theta)^{2}\leq(k+1)/n$. We claim that $\varepsilon^{\ast}\asymp\sqrt{k/n}$, where $\varepsilon^{\ast}$ is as defined in (2.1). As a result, the minimax rate will be $k/n\wedge a_{p}$ if $a_{p}>1/n$, and $a_{p}$ if $a_{p}\leq 1/n$, all up to constants.
We start with the upper bound. In the case $a_{p}\leq 1/n$, note that the minimax rate is trivially bounded by the diameter $4a_{p}$. Suppose instead $a_{p}>1/n.$ Now take $\varepsilon^{2}\geq Ck/n$. Pick any $f_{\theta}\in\mathcal{F}$ and let $M_{f_{\theta}}$ be an $\varepsilon/c$-packing of $\mathcal{F}\cap B(f_{\theta},\varepsilon)$. Then this corresponds to some $\varepsilon/c$-packing $M_{\theta}$ of the set $\Theta\cap B_{2}^{p}(\theta,\varepsilon)$, where we are using the $p$-dimensional Euclidean $l_{2}$-norm in the ball centered at $\theta$. Let $P\in\mathcal{P}_{k}$ solve the outer minimization problem in the Kolmogorov width definition for $\theta$.
Then for any $g_{\theta_{1}},g_{\theta_{2}}\in M_{f_{\theta}}$, we have
$\frac{\varepsilon}{c}$ $\leq\|g_{\theta_{1}}-g_{\theta_{2}}\|_{L_{2}(\mathbb{P}_{X})}=\|\theta_{1}-\theta_{2}\|_{2}\leq\|\theta_{1}-P\theta_{1}-\theta_{2}+P\theta_{2}\|_{2}+\|P\theta_{1}-P\theta_{2}\|_{2}$
$\leq 2d_{k}(\Theta)+\|P\theta_{1}-P\theta_{2}\|_{2}.$
Note that $\varepsilon^{2}\geq Ck/n>\frac{C(k+1)}{2n}\geq\frac{C}{2}\cdot d_{k}(\Theta)^{2},$ which means $2d_{k}(\Theta)\leq\sqrt{\frac{8}{C}}\cdot\varepsilon$ so by taking $C$ sufficiently large, we have
$\|P\theta_{1}-P\theta_{2}\|_{2}$ $\geq\frac{\varepsilon}{c}-2d_{k}(\Theta)>\frac{\varepsilon}{c}-\sqrt{\frac{8}{C}}\cdot\varepsilon=\frac{\varepsilon}{C^{\prime}}$
for some constant $C^{\prime}$. This means we have formed an $\frac{\varepsilon}{C^{\prime}}$-packing set $\{P\theta:\theta\in M_{\theta}\}$ of the $k$-dimensional set $P\Theta$. On the other hand, a $k$-dimensional set contained in a $k$-dimensional sphere has a packing set with log cardinality bounded by $kC^{\prime\prime}$ for some constant $C^{\prime\prime}$. Since $\|P\theta_{1}-P\theta_{2}\|_{2}>0$ for each $\theta_{1},\theta_{2}\in M_{\theta}$, we have $|M_{\theta}|=|\{P\theta:\theta\in M_{\theta}\}|$. Hence the packing set $M_{\theta}$ has log cardinality bounded by $kC^{\prime\prime}$, as does $M_{f_{\theta}}$ by the equivalence of the Euclidean norm and $L_{2}(\mathbb{P}_{X})$ norm.
Hence we have $n\varepsilon^{2}\geq kC$ while $\log M_{\mathcal{F}}^{\operatorname{loc}}(\varepsilon,c)\leq kC^{\prime\prime}$, so taking $C$ large enough, we have $n\varepsilon^{2}>\log M_{\mathcal{F}}^{\operatorname{loc}}(\varepsilon,c).$ But since $n\varepsilon^{2}$ is increasing in $\varepsilon$ while $\log M_{\mathcal{F}}^{\operatorname{loc}}(\varepsilon,c)$ is non-increasing in $\varepsilon$, we must have $\varepsilon^{\ast}\leq\varepsilon$. Since this holds for all $\varepsilon\geq\sqrt{Ck/n}$, it follows that $\varepsilon^{\ast}\leq\sqrt{Ck/n}$. The minimax rate will also be bounded by the diameter, so the minimax rate is therefore upper-bounded up to constants by $\sqrt{k/n}\wedge 4a_{p}$ when $a_{p}>1/n$.
Now we prove the lower bound. We first handle the $a_{p}>1/n$ case. Consider the subset $\Theta^{\prime}=\{\theta\in\mathbb{R}^{p}:\sum_{i=n-k+1}\theta_{i}^{2}/a_{i}\leq 1\}\subseteq\Theta$, i.e., we set the first $n-k$ coordinates of $\theta\in\Theta$ to 00. But as $a_{p-k+1}>k/n$ and $a_{1}\leq a_{2}\leq\dots\leq a_{p}$, we conclude that $\Theta^{\prime}$ and therefore $\Theta$ contains a $k$-dimensional ball of radius $\sqrt{k/n}$. Hence $\log M_{\mathcal{F}}^{\operatorname{loc}}(\varepsilon,c)=\log M_{\Theta}^{\operatorname{loc}}(\varepsilon,c)\geq Ck$ for some constant $C$. Now take $\varepsilon=C^{\prime}\sqrt{k/n}$ for some sufficiently small constant $C^{\prime}$ such that $n\varepsilon^{2}=(C^{\prime})^{2}k\leq Ck=\log M_{\mathcal{F}}^{\operatorname{loc}}(\varepsilon,c)$. Since $\varepsilon^{\ast}$ is the supremum of all such $\varepsilon$ satisfying $n\varepsilon^{2}\leq\log M_{\mathcal{F}}^{\operatorname{loc}}(\varepsilon,c)$, we conclude $\varepsilon^{\ast}\geq\varepsilon\gtrsim\sqrt{k/n}$. Combined with the upper bound, we conclude in the case where $a_{p}>1/n$ that ${\varepsilon^{\ast}}^{2}\asymp k/n$ and thus the minimax rate is $k/n\wedge 4a_{p}$ up to constants.
Lastly, we consider the lower bound for when $a_{p}\leq n^{-1}$, which implies $a_{i}\leq n^{-1}$ for all $i\in[p]$. Now take $\varepsilon=\frac{2\sqrt{a_{p}}}{C\sqrt{n}}$ for some constant $C$. For sufficiently large $C$, we have $\log M_{\mathcal{F}}^{\operatorname{loc}}(\varepsilon,c)>1>\frac{4a_{p}}{C^{2}}=n\cdot\varepsilon^{2}.$ Again by the definition of $\varepsilon^{\ast}$ as a supremum, we have $\varepsilon^{\ast}\geq\varepsilon=\frac{2\sqrt{a_{p}}}{C\sqrt{n}}$, so $\varepsilon^{\ast}\gtrsim\sqrt{a_{p}}$, noting that $\sqrt{a_{p}}$ is the diameter up to constants. So the minimax rate is lower bounded by $a_{p}$ up to constants, and we already know it is upper bounded by it in this case. This completes our proof that the minimax rate is $k/n\wedge 4a_{p}$ if $a_{p}>1/n$, and $a_{p}$ if $a_{p}\leq 1/n$, where $k\in[p]$ is such that $a_{p-k}\leq(k+1)/n$ and $a_{p-k+1}>k/n$.
Now, let us compare our result to that of pathak2023noisy. Define
$d_{n}=\sup_{\Omega\succ 0}\left\{\mathbb{E}\;\mathrm{Tr}\left(\Omega^{-1}+n^{-1}X^{T}X\right)^{-1}:\mathrm{Tr}(K_{c}^{-1/2}\Omega K_{c}^{1/2})\leq n\right\},$
where we recall $K_{c}$ is the diagonal matrix with $(i,i)$-entry $a_{i}$ for each $1\leq i\leq p$. The authors then conclude the minimax rate is of order $d_{n}/n$.
Thus, if $a_{p}\leq 1/n$, we equate the two results and conclude that $a_{p}\asymp d_{n}/n$, or if $a_{p}>1/n$ we conclude $k/n\wedge a_{p}\asymp d_{n}/n$. Although calculating $d_{n}$ from its definition may be challenging, our approach with local metric entropies is able to indirectly recover it.
5.2 Lipschitz Classes
Given any function $f\colon[0,1]\to\mathbb{R}$, we extend $f$ to $(1,\infty)$ by setting $f(x)=f(1)$ for $x>1$. For any such function $f$ and $h>0$, define the function $\Phi_{f,h}\colon[0,1]\to\mathbb{R}$ by $\Phi_{f,h}(x)=f(x+h)-f(x)$. Now for each $\alpha\in(0,1]$ and $\gamma>0$, define
$\mathcal{F}_{\alpha,\gamma}=\{f\colon[0,1]\to\mathbb{R}:\|f\|_{L_{2}(\mathbb{P}_{X})}\leq\gamma,\|\Phi_{f,h}\|_{L_{2}(\mathbb{P}_{X})}\leq\gamma h^{\alpha}\text{ for all }h>0\}.$
Then this is a convex function class with bounded diameter, i.e., $\gamma\leq\mathrm{diam}(\mathcal{F}_{\alpha,\gamma})\leq 2\gamma$.
Then we have $\log M(\varepsilon,\mathcal{F}_{\alpha,\gamma})\asymp\varepsilon^{-1/\alpha}$ by yang1999information and using the inequality
$\log M(\varepsilon/c,\mathcal{F}_{\alpha,\gamma})\geq\log M_{\mathcal{F}_{\alpha,\gamma}}^{\operatorname{loc}}(\varepsilon,c)\geq\log M(\varepsilon/c,\mathcal{F}_{\alpha,\gamma})-\log M(\varepsilon,\mathcal{F}_{\alpha,\gamma}),$
we conclude $\log M_{\mathcal{F}_{\alpha,\beta,\gamma}}^{\operatorname{loc}}(\varepsilon,c)\asymp\varepsilon^{-1/\alpha}$. The relationship $n\varepsilon^{2}\asymp\varepsilon^{-1/\alpha}$ occurs when $\varepsilon\asymp n^{-\frac{\alpha}{2\alpha+1}}$, so we conclude the minimax-rate is determined by $n^{-\frac{\alpha}{2\alpha+1}}\wedge\mathrm{diam}(\mathcal{F}_{\alpha,\gamma})\asymp n^{-\frac{\alpha}{2\alpha+1}}\wedge\gamma.$
5.3 Monotone functions in general dimensions
[…]
$f(x_{1},\dots,x_{i-1},x_{i},x_{i+1},\dots,x_{p})\leq f(x_{1},\dots,x_{i-1},x_{i}+y_{i},x_{i+1},\dots,x_{p})$
for any $(x_{1},\dots,x_{p})\in[0,1]^{p}$, $y_{i}>0$, and $1\leq i\leq p$. We restate a key theorem from gao_Wellner_monotone:
Lemma 5.1 (Theorem 1.1 of gao_Wellner_monotone).
For $p>2$, we have
$\log M(\varepsilon,\mathcal{F}_{p})\asymp\varepsilon^{-2(p-1)}.$
If $p=2$, then
$\varepsilon^{-2}\lesssim\log M(\varepsilon,\mathcal{F}_{p})\lesssim\varepsilon^{-2}(\log 1/\varepsilon)^{2}.$
Assume $p>2$ at first. Recall once more by yang1999information that
$\log M(\varepsilon/c,\mathcal{F}_{p})\geq\log M_{\mathcal{F}_{p}}^{\operatorname{loc}}(\varepsilon,c)\geq\log M(\varepsilon/c,\mathcal{F}_{p})-\log M(\varepsilon,\mathcal{F}_{p}).$
Therefore $\log M_{\mathcal{F}_{p}}^{\operatorname{loc}}(\varepsilon,c)\asymp\varepsilon^{-2(p-1)}$. Next, the relationship $n\varepsilon^{2}\asymp\varepsilon^{-2(p-1)}$ holds when $\varepsilon\asymp n^{-1/(2p)}$. Since the diameter of $\mathcal{F}_{p}$ is 1, we conclude the minimax rate for $p>2$ is given by $n^{-1/p}\wedge 1$ up to constants. Note that this omits the logarithmic factor appearing in samsworth_isotonic. This logarithmic factor could possibly be extraneous, or it is also possible that they have it since they use the LSE rather than our minimax optimal estimator.
Let us next consider the $p=2$ case. Since for any $\varepsilon>0$ satisfying $n\varepsilon^{2}\leq\log M_{\mathcal{F}_{2}}^{\operatorname{loc}}(\varepsilon,c)$ we have $n\varepsilon^{2}\lesssim\varepsilon^{-2}(\log 1/\varepsilon)^{2}$, it follows that $\varepsilon^{\ast}$ as defined in Theorem 2.12 satisfies $n{\varepsilon^{\ast}}^{4}\lesssim(\log 1/\varepsilon^{\ast})^{2}.$ On the other hand, since $\log M_{\mathcal{F}_{2}}^{\operatorname{loc}}(\varepsilon,c)\gtrsim\varepsilon^{-2}$, we have $n{\varepsilon^{\ast}}^{2}\gtrsim{\varepsilon^{\ast}}^{-2}$, i.e., $\varepsilon^{\ast}\gtrsim n^{-1/4}$. Hence
$n^{-1/4}\lesssim\varepsilon^{\ast}\lesssim n^{-1/4}\sqrt{\log(1/\varepsilon^{\ast})}.$
Reusing $n^{-1/4}\lesssim\varepsilon^{\ast}$ note that $\sqrt{\log(1/\varepsilon^{\ast})}\lesssim\sqrt{\log(n^{1/4})}\asymp\sqrt{\log n}.$ Hence $n^{-1/4}\lesssim\varepsilon^{\ast}\lesssim n^{-1/4}\sqrt{\log(n)}.$ The minimax rate therefore satisfies
$n^{-1/2}\wedge 1\lesssim{\varepsilon^{\ast}}^{2}\wedge 1\lesssim(n^{-1/2}\log n)\wedge 1.$
For $p=1$, we know $\log M(\varepsilon,\mathcal{F}_{1})\asymp\varepsilon^{-1}$ by van1996weak. Reusing the result from yang1999information we conclude $\log M_{\mathcal{F}_{2}}^{\operatorname{loc}}(\varepsilon,c)\asymp\varepsilon^{-1}$, and this rate matches $n\varepsilon^{2}$ when $\varepsilon\asymp n^{-1/3}$. The minimax rate is therefore $n^{-2/3}\wedge 1$.